### v2

# Spectral properties of symmetrized AMV operators

Abstract.
The symmetrized Asymptotic Mean Value Laplacian $\tilde{\Delta}$, obtained as limit of approximating operators $\tilde{\Delta}_{r}$, is an extension of the classical Euclidean Laplace operator to the realm of metric measure spaces. We show that, as $r\downarrow 0$, the operators $\tilde{\Delta}_{r}$ eventually admit isolated eigenvalues defined via min-max procedure on any compact uniformly locally doubling metric measure space. Then we prove $L^{2}$ and spectral convergence of $\tilde{\Delta}_{r}$ to the Laplace–Beltrami operator of a compact Riemannian manifold, imposing Neumann conditions when the manifold has a non-empty boundary.
1. Introduction
In the past thirty years, much research has been carried out to extend the classical Euclidean Laplace operator to metric measure spaces: see e.g. [CC00, KMS01, Gig15, AB18]. This paper deals with such an extension, namely the symmetrized Asymptotic Mean Value (AMV) Laplacian, proposed in [MT23], see also [Kok06, MT20, AKS22, AKS23]. The symmetrized AMV Laplacian is set as
$\tilde{\Delta}\coloneqq\lim\limits_{r\downarrow 0}\tilde{\Delta}_{r}$ (1)
where for $\mu$-a.e. $x\in X$,
$\tilde{\Delta}_{r}f(x)\coloneqq\frac{1}{2r^{2}}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)(f(y)-f(x))\mathop{}\!\mathrm{d}\mu(y).$
Here $f$ is a locally integrable function defined on a metric measure space $(X,\mathsf{d},\mu)$. Throughout the paper, $B_{r}(z)$ denotes the metric open ball centered at $z\in X$ with radius $r>0$, the notation $V(z,r)$ stands for $\mu(B_{r}(z))$, and $\fint_{B_{r}(z)}$ is shorthand for $V(z,r)^{-1}\int_{B_{r}(z)}$.
Part of the study on the symmetrized AMV Laplacian consists in finding a relevant meaning to the limit in (1). If this is intended in the $L^{2}$ sense, then the associated spectral convergence can be investigated. This is the goal of the present paper. For any $k\in\mathbb{N}$, set
$\tilde{\lambda}_{k,r}\coloneqq\inf_{V\in\mathcal{G}_{k+1}(L^{2}(X,\mu))}\sup_{f\in V}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}},$ (2)
where $\mathcal{G}_{k+1}(L^{2}(X,\mu))$ is the $(k+1)$-th Grassmannian of $L^{2}(X,\mu)$, and $\tilde{E}_{r}(f)$ is the energy functional naturally associated with $\tilde{\Delta}_{r}$ (Definition 3.4). These form a non-decreasing sequence of non-negative numbers. Our first main result states that these numbers eventually correspond to isolated eigenvalues of $-\tilde{\Delta}_{r}$ when $(X,\mathsf{d},\mu)$ is compact and uniformly locally doubling (Definition 2.1).
Theorem 1.
Let $(X,\mathsf{d},\mu)$ be a compact uniformly locally doubling metric measure space. For any integer $k\geq 2$, there exists $r_{k}>0$ such that for any $r\in(0,r_{k})$, the operator $-\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues
$0=\lambda_{0}(-\tilde{\Delta}_{r})<\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$
such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$.
Our second main result deals with a smooth manifold $M$ endowed with a smooth Riemannian metric $g$. We write $\Delta_{g}$ for the (negative) Laplace–Beltrami operator of $(M,g)$. We let $m\geq 2$ be the dimension of $M$, and we set
$C_{m}\coloneqq\frac{1}{2}\fint_{\mathbb{B}^{m}_{1}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\xi=\frac{1}{2(m+2)}$ (3)
where $\mathbb{B}^{m}_{1}(0)$ is the unit Euclidean ball of $\mathbb{R}^{m}$. In this context, it follows from the equality between symmetrized and non-symmetrized AMV Laplacian and a simple calculation in normal coordinates that
$\tilde{\Delta}_{r}f(x)\stackrel{{\scriptstyle r\downarrow 0}}{{\to}}C_{m}\Delta_{g}f(x)$ (4)
for any $f\in\mathcal{C}^{2}(M)$ and any interior point $x\in M$, see [MT20, MT23] — the convergence is even locally uniform in the interior of $M$, see [AKS22]. We refer to [MT20, MT23, AKS22, AKS23] for related pointwise results in various settings like Carnot groups or Alexandrov spaces.
In this paper, we are interested in the $L^{2}$ version of (4) with a particular interest in the case where $M$ admits a non-empty boundary $\partial M\neq 0$. In this case, we write $\partial_{\nu}f\in\mathcal{C}^{\infty}(\partial M)$ for the normal derivative of a smooth function $f:M\to\mathbb{R}$, and we define
$\mathcal{C}^{\infty}_{\nu}(M)\coloneqq\{f\in\mathcal{C}^{\infty}(M)\,:\,\partial_{\nu}f=0\}.$ (5)
We see $(M,g)$ as a metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ where $\mathsf{d}_{g}$ and $\mathrm{vol}_{g}$ are the Riemannian distance and volume measure on $M$ associated with $g$. Then our statement reads as follows.
Theorem 2.
Let $(M^{m},g)$ be a compact, connected, smooth Riemannian manifold with a non-empty (resp. empty) boundary $\partial M$. Then for any $f\in\mathcal{C}_{\nu}^{\infty}(M)$ (resp. $\mathcal{C}^{\infty}(M)$), as $r\downarrow 0$,
$\tilde{\Delta}_{r}f\stackrel{{\scriptstyle L^{2}}}{{\longrightarrow}}C_{m}\Delta_{g}f.$
We point out that the boundaryless version of this result is rather easy to obtain, while a non-empty boundary is quite tricky to handle. The Neumann condition in the latter case is crucial to ensure convergence: indeed, the sequence $\tilde{\Delta}_{r}f$ may blow-up if this is not imposed.
After the previous $L^{2}$-convergence result, we address the question of spectral convergence, that is to say, the convergence of the associated eigenvalues and eigenfunctions. In this regard we show that, for any $k\in\mathbb{N}$, the function $r\mapsto\tilde{\lambda}_{k,r}$ is bounded in a neighborhood of 00, as proved in the course of Theorem 1. This ensures that the $k$-th lowest eigenvalue of the operator $-\tilde{\Delta}_{r}$, which we denote $\lambda_{k}(-\tilde{\Delta}_{r})$, exists for small enough $r$, and that it coincides with $\tilde{\lambda}_{k,r}$. Let $f_{k,r}$ be an $L^{2}$-normalized eigenfunction of $-\tilde{\Delta}_{r}$ associated with $\lambda_{k}(-\tilde{\Delta}_{r})$. Recall that if $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$), a Laplace (resp. Neumann) eigenvalue of $(M,g)$ is a number $\mu\geq 0$ for which there exists an associated eigenfuction $f\in\mathcal{C}^{\infty}(M)$ (resp. $\mathcal{C}_{\nu}^{\infty}(M)$) of $-\Delta_{g}$, i.e. $-\mu f=\Delta_{g}f$.
Theorem 3.
Let $(M^{m},g)$ be a compact, connected, smooth Riemannian manifold. Assume that $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$). For $k\in\mathbb{N}$, let $\mu_{k}$ be the $k$-th lowest Laplace (resp. Neumann) eigenvalue of $\Delta_{g}$. For any $(r_{n})\subset(0,+\infty)$ such that $r_{n}\to 0$, there exists an $L^{2}$-normalized Laplace (resp. Neumann) eigenfunction $f\in\mathcal{C}^{\infty}(M)$ (resp. $\mathcal{C}_{\nu}^{\infty}(M)$) associated with $\mu_{k}$ such that, up to a subsequence,
$\begin{cases}\lambda_{k}(-\tilde{\Delta}_{r_{n}})\to C_{m}\,\mu_{k},\\ f_{k,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\longrightarrow}}f.\end{cases}$ (6)
We point out that the question of spectral convergence for the Gaussian approximation of the Laplace–Beltrami operator of a compact Euclidean submanifold with boundary was raised in [BQWZ12]. This has been one motivation for the present work: to study this convergence with the intrinsic approximation provided by the symmetrized AMV operators $\tilde{\Delta}_{r}$ instead of the extrinsic Gaussian one.
Acknowledgments.
Both authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They thank the anonymous reviewers for their constructive suggestions, and Dmitri Pavlov for clarifying a point related to Lemma 2.1.
2. Averaging-like operators
In this section, we consider a fixed metric measure space, that is to say, a triple $(X,\mathsf{d},\mu)$ where $(X,\mathsf{d})$ is a metric space and $\mu$ is a fully supported regular Borel measure on $(X,\mathsf{d})$ such that
$V(x,r)\coloneqq\mu(B_{r}(x))<+\infty$
for any $x\in X$ and $r>0$, where $B_{r}(x)$ denotes the open ball $\{y\in X:\mathsf{d}(x,y)<r\}$. Notice that for any $x\in X$ and $r>0$,
$V(x,r)>0,$
because $\mu$ is fully supported. Moreover, if $X$ is compact, then
$\mu(X)<+\infty$ (7)
since $\mu$ is finite on any ball of radius the diameter of $X$. We set
$0\leq m(r)\coloneqq\inf_{x\in X}V(x,r)\leq M(r)\coloneqq\sup_{x\in X}V(x,r)\leq+\infty.$
Note that our assumptions yield the following preliminary result.
Lemma 2.1.
$L^{2}(X,\mu)$ is separable.
Proof.
We start by proving that $(X,\mathsf{d})$ is a second countable space. Fix $o\in X$. Given $\epsilon>0$ and $N\in\mathbb{N}$ positive, consider the value given by
$\alpha_{\epsilon,N}=\sup\{\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(x_{n})):\{x_{n}\}_{n\in\mathbb{N}}\subset X\},$ (8)
where $\left.\mu\right|_{B_{N}(o)}(\cdot)\coloneqq\mu(\cdot\cap B_{N}(o))$. First we show this supremum is attained. Consider $\delta_{k}\rightarrow 0$ and let $\{x_{n}^{k}\}_{n\in\mathbb{N}}\subset X$ such that
$\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(x_{n}^{k}))>\alpha-\delta_{k}.$ (9)
Taking
$\{y_{n}\}_{n\in\mathbb{N}}=\cup_{k}\{x_{n}^{k}\}_{n\in\mathbb{N}}$ (10)
we have that
$\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(y_{n}))=\alpha_{\epsilon,N}.$ (11)
Now we prove $\alpha_{\epsilon,N}=\mu(B_{N}(z))<\infty$. If $\alpha_{\epsilon,N}<\mu(B_{N}(z))$, then $\mu(B_{N}(z)\backslash\cup_{n}B_{\epsilon}(y_{n}))>0$, where $\{y_{n}\}_{n\in\mathbb{N}}$ is a maximizer of (8). Since the measure is inner regular, there must exist some compact $K\subset B_{N}(z)\backslash\cup_{n}B_{\epsilon}(y_{n})$ such that
$\mu(K)>0.$ (12)
Since we can cover $K$ by a finite number of balls $B_{\epsilon}(z_{k})$, there must exist some $z=z_{k}$ such that
$\mu(B_{\epsilon}(z)\cap K)>0.$ (13)
We then have
$\left.\mu\right|_{B_{N}(o)}\left(\cup_{n}B_{\epsilon}(y_{n})\cup B_{\epsilon}(z)\right)$ $\geq\left.\mu\right|_{B_{N}(o)}\left(\cup_{n}B_{\epsilon}(y_{n})\cup\left(B_{\epsilon}(z)\cap K\right)\right)$ (14)
$=\mu\left(\cup_{n}B_{\epsilon}(y_{n})\right)+\mu\left(B_{\epsilon}(z)\cap K\right)>\alpha_{N,\epsilon}.$ (15)
And so $\{y_{n}\}_{n\in\mathbb{N}}\cup\{z\}$ contradicts the maximality of $\{y_{n}\}_{n\in\mathbb{N}}$. This shows that $\alpha_{\epsilon,N}=\mu(B_{N}(o))$. Since $\cup_{n}B_{\epsilon}(y_{n})\cap B_{N}(o)$ has full measure in the support $B_{N}(o)$ of $\mu|_{B_{N}(o)}$, it is a dense subset of $B_{N}(o)$. This implies that $\cup_{n}B_{2\epsilon}(y_{n})$ is a countable cover of $B_{N}(o)$. To build a countable basis of $X$, consider a sequence $\delta_{k}\rightarrow 0$. For any $k$, take $\cup_{n\in\mathbb{N}}B_{\delta_{k}}(y_{n}^{k,N})$ a countable cover of $B_{N}(z)$. Then the set given by
$\mathcal{B}=\cup_{k,N\in\mathbb{N}}\{B_{\delta_{k}}(y_{n}^{k,N})\}_{n\in\mathbb{N}}$ (16)
is a countable basis of $X$. Given this basis $\mathcal{B}$ we can create a new basis given by the finite union of elements of $\mathcal{B}$, and we call this new basis $\mathcal{B}^{\prime}$ which will also be countable. In particular we have that given some open set $V\subset X$ we can find a sequence $V_{n}$ such that
$V_{n}\subset V_{n+1},\quad\quad\cup_{n}V_{n}=V.$ (17)
This is the case since $\mathcal{B}$ is a countable basis, we can find elements $B_{k}\in\mathcal{B}$ such that
$\cup_{k}B_{k}=V.$ (18)
We conclude by taking $V_{n}=\cup_{k=1}^{n}B_{k}\in\mathcal{B}^{\prime}$. To construct our dense subset of $L^{2}(X,\mu)$ we take finite sums with rational coefficients of the characteristic functions $\chi_{V}$, with $V\in\mathcal{B}^{\prime}$. To show that this is dense in $L^{2}(X,\mu)$ we only need to show that we can approximate arbitrarily well simple functions $\chi_{U}$ where $U\subset X$ is open and $\mu(U)<\infty$ since the measure $\mu$ is outer regular. Given such an open set $U$, take $\chi_{U_{n}}$ where $U_{n}\in\mathcal{B}^{\prime}$ and $U_{n}\subset U_{n+1}$ and $\cup_{n}U_{n}=U$. Then we have by dominated convergence
$\chi_{U_{n}}\rightarrow_{L^{2}(X,\mu)}\chi_{U},$ (19)
concluding the proof. ∎
2.1. Averaging operator
For any $x,y\in X$ and $r>0$, set
$a_{r}(x,y)\coloneqq\frac{1_{B_{r}(x)}(y)}{V(x,r)}\,\cdot$
Consider $u\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $u$ is $\mu$-integrable on $B_{r}(x)$, set
$A_{r}u(x)\coloneqq\fint_{B_{r}(x)}u\mathop{}\!\mathrm{d}\mu=\int_{X}a_{r}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$
Notice that, since $u$ is locally integrable, for any $x\in X$ there exists $r_{x}>0$ such that $A_{r_{x}}u(x)$ is well-defined. However, there may be no uniform $r>0$ for which the integral $A_{r}u(x)$ is well-defined for every $x\in X$.
In the next lemma, we provide a mild condition under which $A_{r}$ defines a bounded linear operator between Lebesgue spaces.
Lemma 2.2.
Assume that there exists $r>0$ such that
$V(\cdot,r)^{-1}\in L^{1}(X,\mu).$ ($\mathrm{A}_{r}$)
Then for any $p\in[1,+\infty]$ the linear operator $A_{r}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}\|_{p\to p}\leq\|V(\cdot,r)^{-1}\|_{1}^{1/p}\cdot$
Moreover, $A_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
Proof.
The case $p=+\infty$ is obvious, hence we assume $p<+\infty$. Let $u\in L^{p}(X,\mu)$. By Jensen’s inequality, for any $x\in X$,
$|A_{r}u(x)|^{p}\leq\left(\fint_{B_{r}(x)}|u|^{p}\mathop{}\!\mathrm{d}\mu\right).$
Thus
$\|A_{r}u\|_{p}^{p}$ $\leq\int_{X}\frac{1}{V(x,r)}\int_{B_{r}(x)}|u(y)|^{p}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$\leq\left(\int_{X}\frac{1}{V(x,r)}\mathop{}\!\mathrm{d}\mu(x)\right)\|u\|_{p}^{p}.$
Since
$\int_{X}\int_{X}a_{r}^{2}(x,y)\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)=\int_{X}\frac{1}{V(x,r)}\fint_{B_{r}(x)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)=\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}$
we obtain that $A_{r}$ is a Hilbert-Schmidt integral operator acting on the separable space $L^{2}(X,\mu)$ (recall Lemma 2.1); in particular, it is compact [RS80, Section IV.6]. ∎
Lemma 2.3.
Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{A}_{r}$) holds for any $r>0$.
Proof.
Consider $r>0$ and a finite cover $\{B_{r/2}(x_{i})\}$ of $X$. For any $x\in X$, there exists $i$ such that $x\in B_{r/2}(x_{i})$. Then $B_{r}(x)$ contains $B_{r/2}(x_{i})$ so $V(x,r)\geq V(x_{i},r/2)\geq\min_{j}V(x_{j},r/2)>0$. Thus
$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{\min_{j}V(x_{j},r/2)}<+\infty.$
Remark 2.4.
Notice that if $\mu(X)=+\infty$ and $M(r)<+\infty$ then ($\mathrm{A}_{r}$) cannot hold :
$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\geq\frac{\mu(X)}{M(r)}=+\infty.$
This happens, for instance, on $\mathbb{R}^{n}$ endowed with the Euclidean distance and the Lebesgue measure. More generally, this property holds on any locally compact topological group endowed with a left-invariant metric compatible with the Haar measure and with infinite volume (see [Str74, Lemma 1] for more details about these spaces). Moreover, if $\mu(X)<\infty$ and $m(r)>0$, then ($\mathrm{A}_{r}$) always holds :
$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{m(r)}<+\infty.$
In this regard, observe that if $X$ is not totally bounded, then there exist $r>0$ small enough and a countable family of disjoint balls $\{B_{r}(x_{i})\}$ in $X$, so that $\mu(X)<\infty$ and $m(r)>0$ cannot hold simultaneously:
$\mu(X)\geq\sum_{i}V(x_{i},r).$
In the next statement, we provide an alternative way to prove the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to itself. This goes through the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to the space of continuous functions $\mathcal{C}(X)$ which we obtain for $X$ compact and satisfying the following condition:
$\sup_{x\in X}\mu(S_{r}(x))=0$ ($\mathrm{S}_{r}$)
where $S_{r}(x)\coloneqq\{y\in X:\mathsf{d}(x,y)=r\}$.
Lemma 2.5.
Assume that $(X,\mathsf{d},\mu)$ is compact and satisfies ($\mathrm{S}_{r}$) for some $r>0$. Then $A_{r}:L^{2}(X,\mu)\to\mathcal{C}(X)$ is compact and satisfies
$\|A_{r}\|_{2\to\infty}\leq\frac{1}{m(r)^{1/2}}\,\cdot$ (20)
Proof.
[…]
$\leq\mu(B_{r+d(x,z)}(x)-B_{r-d(x,z)}(x))^{\frac{1}{2}},$ (22)
and $\mu(B_{r+d(x,z)}(x)-B_{r-d(x,z)}(x))\to 0$ as $\mathsf{d}(x,z)\to 0$ due to ($\mathrm{S}_{r}$). Moreover, the bound (20) is obtained via Hölder’s inequality: for any $x\in X$,
$|A_{r}u(x)|\leq\left(\fint_{B_{r}(x)}u^{2}\mathop{}\!\mathrm{d}\mu\right)^{1/2}\leq\frac{1}{m(r)^{1/2}}\,\cdot$
To prove compactness, consider $\{f_{n}\}\subset L^{2}(X,\mu)$ such that $\sup_{n}\|f_{n}\|_{2}\leq 1$. Uniform boundedness of $\{A_{r}(f_{n})\}$ follows from (20), and equicontinuity can be obtained by using the inequality (22) applied to the sequence. By the Ascoli-Arzelà Theorem, we can extract from $\{A_{r}(f_{n})\}$ a subsequence which converges in $\mathcal{C}(X)$, concluding the proof. ∎
2.2. Adjoint
For any $x,y\in X$ and $r>0$, set
$a^{*}_{r}(x,y)\coloneqq a_{r}(y,x)=\frac{1_{B_{r}(x)}(y)}{V(y,r)}\,\cdot$
Consider $u\in L^{0}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $v$ is $\mu$-integrable on $B_{r}(x)$, set
$A^{*}_{r}u(x)\coloneqq\int_{B_{r}(x)}\frac{u(y)\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}=\int_{X}a_{r}^{*}(x,y)\mathop{}\!\mathrm{d}\mu(y).$
Notice that, just like $A_{r}u(x)$ in the previous subsection, $A^{*}_{r}u(x)$ may not make sense uniformly with respect to $x\in X$.
Let us begin with a simple observation.
Lemma 2.6.
The operator $A_{r}^{*}:L^{1}(X,\mu)\to L^{1}(X,\mu)$ is a contraction.
Proof.
For any $u\in L^{1}(X,\mu)$,
$\int_{X}|A_{r}^{*}u(x)|\mathop{}\!\mathrm{d}\mu(x)$ $\leq\int_{X}\int_{B_{r}(x)}\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\int_{X}1_{B_{r}(x)}(y)\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\left(\int_{X}1_{B_{r}(y)}(x)\mathop{}\!\mathrm{d}\mu(x)\right)\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)=\int_{X}|u(y)|\mathop{}\!\mathrm{d}\mu(y),$
where we used the Fubini–Tonelli theorem to get the penultimate equality. ∎
The next lemma gives a mild sufficient condition ensuring the well-posedeness and boundedness of the linear operator $A_{r}^{*}$ between $L^{p}$ spaces for $p>1$.
Lemma 2.7.
Assume that there exists $r>0$ such that
$\left.\begin{array}[]{l}V(\cdot,r)^{-1}\in L^{\infty}(B_{r}(x),\mu)\qquad\forall x\in X,\\ \|A_{r}^{*}1\|_{\infty}<+\infty.\end{array}\right\}$ ($\mathrm{A}_{r}^{*}$)
Then for any $p\in[1,+\infty]$, the linear operator $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}^{*}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{(p-1)/p}.$
Moreover, if ($\mathrm{A}_{r}$) additionally holds, then $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is the adjoint of $A_{r}:L^{q}(X,\mu)\to L^{q}(X,\mu)$ for $q\in[1,+\infty]$ such that $11/p+1/q=$ 1, and the operator $A_{r}^{*}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
Proof.
For the proof of the first assertion, consider $u\in L^{\infty}(X,\mu)$. Then for any $x\in X$,
$|A_{r}^{*}u|(x)$ $\leq\int_{B_{r}(x)}\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq\|u\|_{\infty}\int_{B_{r}(x)}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\leq\|A_{r}^{*}1\|_{\infty}\|u\|_{\infty}.$
Thus $A_{r}^{*}:L^{\infty}(X,\mu)\to L^{\infty}(X,\mu)$ is bounded with $\|A_{r}^{*}\|_{\infty\to\infty}\leq\|A_{r}^{*}1\|_{\infty}$. The conclusion for $p\in(1,+\infty)$ follows from the Riesz-Thorin theorem and Lemma 2.6.
Assume now that ($\mathrm{A}_{r}$) holds, and consider $u\in L^{p}(X,\mu)$ and $v\in L^{q}(X,\mu)$. Then
$\int_{X}A_{r}^{*}u(x)\,v(x)\mathop{}\!\mathrm{d}\mu(x)$ $=\int_{X}\int_{X}1_{B_{r}(x)}(y)\,\frac{u(y)}{V(y,r)}\,v(x)\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}\frac{u(y)}{V(y,r)}\int_{X}1_{B_{r}(x)}(y)\,\,v(x)\mathop{}\!\mathrm{d}\mu(x)\mathop{}\!\mathrm{d}\mu(y)$
$=\int_{X}u(y)A_{r}v(y)\mathop{}\!\mathrm{d}\mu(y)$
where we used the Fubini–Tonelli to get the second equality, and the symmetry of the distance function through the equality $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ to get the last one. As for the compactness of $A_{r}^{*}$, the result is a direct consequence of the Schauder theorem for compact operators which can be applied thanks to Lemma 2.2. ∎
Lemma 2.8.
Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{A}_{r}^{*}$) holds for any $r>0$.
Proof.
As seen in the proof of Lemma 2.3, total boundedness implies $m(r)>0$ for any $r>0$ hence $V(\cdot,r)^{-1}\in L^{\infty}(B_{r}(x),\mu)$ for any $x\in X$. Total boundedness also implies that $\mu(X)<+\infty$, thus $\|A_{r}^{*}1\|_{\infty}\leq\mu(X)/m(r)<+\infty$. ∎
Remark 2.9.
Note that if $m(r)>0$ and $M(r)<+\infty$, then ($\mathrm{A}_{r}^{*}$) always holds with
$\|A_{r}^{*}1\|_{\infty}\leq\frac{M(r)}{m(r)}\,\cdot$
This notably happens on locally compact topological groups endowed with a left-invariant distance and their Haar measure, compare with Remark 2.4. Furthermore, observe that $A_{r}^{*}1\in L^{\infty}(X,\mu)$ can be seen as a weak variant of the comparability conditions introduced in [Ald18, MT23].
2.3. Symmetrization
For any $x,y\in X$ and $r>0$, set
$\tilde{a}_{r}(x,y)$ $=\frac{1}{2}\left(a_{r}(x,y)+a^{*}_{r}(x,y)\right)$
$=\frac{1}{2}\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)1_{B_{r}(x)}(y)$
Consider $u\in L^{1}_{\text{loc}}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $u$ and $v$ are $\mu$-integrable on $B_{r}(x)$, set
$\tilde{A}_{r}u(x)\coloneqq\frac{1}{2}(A_{r}u(x)+A_{r}^{*}u(x))=\int_{X}\tilde{a}_{r}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$ (23)
Then the next lemma is an obvious consequence of Lemma 2.2 and Lemma 2.7.
Corollary 2.10.
Assume that there exists $r>0$ such that ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) hold. Then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is a compact self-adjoint operator such that
$\|\tilde{A}_{r}\|_{2\to 2}\leq\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}$
2.4. Doubling condition
Let us recall the classical doubling condition and its uniform local variant, see e.g. [HKST15] for more details.
Definition 2.1.
The space $(X,\mathsf{d},\mu)$ is called doubling if there exists $C>0$ such that for any $x\in X$ and $r>0$,
$V(x,2r)\leq CV(x,r).$ (24)
It is called uniformly uniformly locally doubling if there exist $C,r_{0}>0$ such that (24) holds for any $x\in X$ and $r\in(0,r_{0})$.
It is worth pointing out that the celebrated Bishop–Gromov theorem, see e.g. [Cha95, Theorem III.4.5], implies that any complete Riemannian manifold with a uniform lower bound on the Ricci curvature is uniformly uniformly locally doubling. This is also true for metric spaces with generalized sectional curvature bounded from below in the sense of Alexandrov [BBI01, Theorem 10.6.6] and $\mathrm{CD}(K,N)$ metric measure space [Vil09, Corollary 30.14].
3. Symmetrized AMV operators
In this section, we provide our working definition of the symmetrized AMV $r$-Laplace operator $\tilde{\Delta}_{r}$ and we derive several spectral properties in a general setting.
3.1. Definitions
For this subsection, we consider a metric measure space $(X,\mathsf{d},\mu)$ satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$.
Definition 3.1.
The symmetrized AMV $r$-Laplace operator of $(X,\mathsf{d},\mu)$ is
$\tilde{\Delta}_{r}\coloneqq\frac{1}{r^{2}}(\tilde{A}_{r}-[\tilde{A}_{r}1]\mathrm{I})$
where we recall that $\tilde{A}_{r}$ is defined in (23).
Remark 3.1.
We may use the notation $\tilde{\Delta}_{r,\mathfrak{X}}$ to specify that we work on the metric measure space $\mathfrak{X}=(X,\mathsf{d},\mu)$.
Lemma 3.2.
$\tilde{\Delta}_{r}$ is a bounded, self-adjoint operator acting on $L^{2}(X,\mu)$ with
$\|\tilde{\Delta}_{r}\|_{2\to 2}\leq\frac{1}{r^{2}}\left(\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}+c(r)+1\right).$ (25)
Proof.
The self-adjointness of $\tilde{\Delta}_{r}$ is obvious because $\tilde{A}_{r}$ and $[\tilde{A}_{r}1]\mathrm{I}$ are self-adjoint too. The boundedness is a consequence of Lemma 2.10. Indeed,
$\|\tilde{\Delta}_{r}\|_{2\to 2}$ $\leq\frac{1}{r^{2}}\left(\|\tilde{A}_{r}\|_{2\to 2}+\|[\tilde{A}_{r}1]I\|_{2\to 2}\right)$
$\leq\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\|\tilde{A}_{r}1\|_{\infty}\underbrace{\|I\|_{2\to 2}}_{=1}\right)$
$\leq\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\frac{\|A_{r}1\|_{\infty}+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
$=\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\frac{1+c(r)}{2}\right)$
hence $\tilde{\Delta}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is bounded and (25) holds. ∎
Definition 3.2.
The energy functional $\tilde{E}_{r}$ of $(X,\mathsf{d},\mu)$ is the quadratic form on $L^{2}(X,\mu)$ defined by
$\tilde{E}_{r}(f)\coloneqq\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\left(\frac{f(x)-f(y)}{r}\right)^{2}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
The associated bilinear form, which we still denote by $\tilde{E}_{r}$, is given by
$\tilde{E}_{r}(f,\psi)$ $\coloneqq\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)$
×$\qquad\qquad\qquad\qquad\times\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
Remark 3.3.
We may also use the notation $\tilde{E}_{r,\mathfrak{X}}$ to specify the metric measure space $\mathfrak{X}=(X,\mathsf{d},\mu)$.
The next lemma goes back to [AKS23, Lemma 3.1]. We provide a quick proof for completeness.
Lemma 3.4.
For any $f,\psi\in L^{2}(X,\mu)$,
$\tilde{E}_{r}(f,\psi)=\langle-\tilde{\Delta}_{r}f,\psi\rangle_{L^{2}}.$ (26)
Proof.
Note that
$\tilde{E}_{r}(f,\psi)$ $=\frac{1}{4}\int_{X}\int_{B_{r}(x)}\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\frac{(f(x)-f(y))\psi(x)}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$-\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\frac{(f(x)-f(y))\psi(y)}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
Using $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ and then the Fubini theorem, we can rewrite the second term as the opposite of the first one, so that we eventually get (26). ∎
Remark 3.5.
Observe that (26) implies that $-\tilde{\Delta}_{r}$ is a non-negative operator, since for any $f\in L^{2}(X,\mu)$,
$\langle-\tilde{\Delta}_{r}f,f\rangle_{L^{2}}=\tilde{E}_{r}(f)\geq 0$
Let us recall the definition of spectrum.
Definition 3.3.
We let $\sigma(-\tilde{\Delta}_{r})$ denote the spectrum of $-\tilde{\Delta}_{r}$, that is to say, the set of elements $\lambda\in\mathbb{C}^{*}$ such that $-\tilde{\Delta}_{r}-\lambda\mathrm{I}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is not a bijection.
It is well-known from classical functional analysis that the spectrum $\sigma(T)$ of a bounded operator $T$ acting on a Banach space $E$ can be decomposed as
$\sigma(T)=\sigma_{\text{p}}(T)\cup\sigma_{\text{c}}(T)\cup\sigma_{\text{a}}(T)$
where:
  • $\sigma_{\text{p}}(T)$ is the point spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ such that $(T-\lambda\mathrm{I})f=0$ for some non-zero $f\in E$, in which case $\lambda$ is called an eigenvalue and $f$ an eigenvector of $T$,
  • $\sigma_{\color[rgb]{.75,0,.25}\text{c}}(T)$ is the compression spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ whose conjugate $\bar{\lambda}$ is an eigenvalue of the adjoint $T^{*}$,
  • $\sigma_{\text{a}}(T)$ is the approximate point spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ for which there exists $(f_{n})\subset E$ with $\|f_{n}\|=1$ for any $n$ such that $\|(T-\lambda\mathrm{I})f_{n}\|\to 0$.
Since $-\tilde{\Delta}_{r}$ is self-adjoint and non-negative, we know that
$\sigma(-\tilde{\Delta}_{r})\subset[0,+\infty].$
This implies that $\sigma_{\text{p}}(-\tilde{\Delta}_{r})=\sigma_{\text{c}}(-\tilde{\Delta}_{r})$, so that
$\sigma(-\tilde{\Delta}_{r})=\sigma_{\text{p}}(-\tilde{\Delta}_{r})\cup\sigma_{\text{a}}(-\tilde{\Delta}_{r}).$ (27)
Definition 3.4.
For any $k\in\mathbb{N}$, we define
$\tilde{\lambda}_{k,r}\coloneqq\inf_{V\in\mathcal{G}_{k+1}(L^{2}(X,\mu))}\sup_{f\in V}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}},$ (28)
where $\mathcal{G}_{k+1}(L^{2}(X,\mu))$ is the $(k+1)$-th Grassmannian of $L^{2}(X,\mu)$.
Remark 3.6.
Let $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ denote the essential spectrum of $-\tilde{\Delta}_{r}$, i.e. the closed subset of $\sigma(-\tilde{\Delta}_{r})$ made of those $\lambda$ such that $-\tilde{\Delta}_{r}-\lambda\mathrm{I}$ is not a Fredholm operator. Since $-\tilde{\Delta}_{r}$ is self-adjoint, the Fischer–Polyà minimum-maximum principle (see e.g. [WS72, p.12]) asserts that if there exists a positive integer $N$ such that $\tilde{\lambda}_{N,r}<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ then $-\tilde{\Delta}_{r}$ admits $N+1$ isolated eigenvalues
$\lambda_{0}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{N}(-\tilde{\Delta}_{r})<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$
such that for any $k\in\{0,\ldots,N\}$,
$\tilde{\lambda}_{k,r}=\lambda_{k}(-\tilde{\Delta}_{r}).$
3.2. Spectral properties
Our first spectral result on $\tilde{\Delta}_{r}$ is the following.
Proposition 3.7.
Let $(X,\mathsf{d},\mu)$ be a metric measure space satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$. Assume that $\lambda\in\sigma(-\tilde{\Delta}_{r})$ satisfies
$\lambda<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}.$ (29)
Then $\lambda$ is an isolated eigenvalue of $-\tilde{\Delta}_{r}$ which does not belong to $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$.
Proof.
Let us first show that $\lambda$ is an eigenvalue, that is to say, that $\lambda$ belongs to the point spectrum. According to (27), it is enough to show that if $\lambda$ is in the approximate point spectrum, then it is in the point spectrum. If this is the case, then there exists a sequence $(f_{n})\in L^{2}(X,\mu)$ such that $\|f_{n}\|_{L^{2}(X)}=1$ and
$\|-\tilde{\Delta}_{r}f_{n}-\lambda f_{n}\|_{L^{2}(X)}=\left\|\left(\frac{[\tilde{A}_{r}1]}{r^{2}}-\lambda\right)f_{n}-\frac{\tilde{A}_{r}f_{n}}{r^{2}}\right\|_{L^{2}(X)}\to 0.$ (30)
Since $\tilde{A}_{r}$ is compact, we have that $\tilde{A}_{r}f_{n}$ converges up to a subsequence, and as such by equation (30) so does $\left([\tilde{A}_{r}1]-r^{2}\lambda\right)f_{n}$, with limit $g\in L^{2}(X)$. Consider $\delta>0$ such that $0\leq\lambda+\delta/r^{2}<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}$ and define
$b_{r}(x):=[\tilde{A}_{r}1]-r^{2}\lambda\geq\delta.$ (31)
Thus we have that $f:=g/b_{r}\in L^{2}(X,\mu)$, and so
$\delta\|f_{n}-\frac{g}{b_{r}}\|_{L^{2}(X)}\leq\|b_{r}f_{n}-g\|_{L^{2}(X)}\rightarrow 0.$ (32)
Thus $f_{n}$ converges in $L^{2}(X,\mu)$ to the limit function $f$. Using continuity of $-\tilde{\Delta}_{r}$ we conclude that
$-\tilde{\Delta}_{r}f=\lambda f,$ (33)
and thus $\lambda$ is in the point spectrum.
To prove that $\lambda$ is an isolated point, we suppose by contradiction that there exists an infinite sequence $(\lambda_{n})\subset\sigma(-\tilde{\Delta}_{r})$ of distinct values such that $\lambda_{n}\rightarrow\lambda$. Then there exists $\delta>0$ such that for any high enough $n$,
$\lambda_{n}+\delta/r^{2}<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}},\quad\quad\lambda_{n}\rightarrow\lambda.$ (34)
From the previous paragraph, we know that $\lambda_{n}$ is in the point spectrum, thus there exists $f_{n}\in L^{2}(X,\mu)$ satisfying $\|f_{n}\|_{L^{2}(X)}=1$, such that
$-\tilde{\Delta}_{r}f_{n}=\lambda_{n}f_{n}.$ (35)
This can be written as
$-\frac{\tilde{A}_{r}}{r^{2}}f_{n}=(\lambda_{n}-\frac{[\tilde{A}_{r}1]}{r^{2}})f_{n}.$ (36)
Using compactness of $\tilde{A}_{r}$, we know that $\tilde{A}_{r}f_{n}$ converges up to a subsequence. This implies that $(r^{2}\lambda_{n}-[\tilde{A}_{r}1])f_{n}$ converges up to a subsequence to some $g\in L^{2}(X,\mu)$. Define
$b_{r,n}(x):=[\tilde{A}_{r}1]-r^{2}\lambda_{n}.$ (37)
With this we have that
$\delta\left\|f_{n}-\frac{g}{b_{r}}\right\|_{L^{2}(X)}$ $\leq\delta\left(\left\|f_{n}-\frac{g}{b_{r,n}}\right\|_{L^{2}(X)}+\left\|\frac{g}{b_{r,n}}-\frac{g}{b_{r}}\right\|_{L^{2}(X)}\right)$ (38)
$\leq\|b_{r,n}f_{n}-g\|_{L^{2}(X)}+\delta\left\|\frac{g}{b_{r,n}}-\frac{g}{b_{r}}\right\|_{L^{2}(X)}.$ (39)
We have that $\|b_{r,n}f_{n}-g\|_{L^{2}(X)}\rightarrow 0$ and also $\|\frac{g}{b_{r,n}}-\frac{g}{b_{r}}\|_{L^{2}(X)}\rightarrow 0$ since $0<\delta\leq b_{r,n},b_{r}$ and $\lambda_{n}\rightarrow\lambda$. Thus $f_{n}$ converges. However since all the eigenvalues are different, we know that $\langle f_{n},f_{j}\rangle=\delta_{n,j}$, and so the sequence cannot converge up to a subsequence, achieving contradiction. This shows that $\lambda$ is an isolated point of $\sigma(-\tilde{\Delta}_{r})$ finishing the first part of the proof.
Let us now prove that $-\tilde{\Delta}_{r}-\lambda I$ is a Fredholm operator.
To show that $\ker(-\tilde{\Delta}_{r}-\lambda I)$ is finite dimensional we proceed by contradiction. Assume that there exists an infinite sequence $(f_{n})\subset\ker(-\tilde{\Delta}_{r}-\lambda I)$ such that $\langle f_{n},f_{j}\rangle=\delta_{n,j}$. Thus
$-\frac{\tilde{A}_{r}f_{n}}{r^{2}}=\left(\lambda-\frac{[\tilde{A}_{r}1]}{r^{2}}\right)f_{n}.$ (40)
Similar to before we can use compactness of $\tilde{A}_{r}$ and the condition on $\lambda$ to conclude that $f_{n}$ converges in $L^{2}(X,\mu)$ up to a subsequence. However, this is prevented by $\langle f_{n},f_{j}\rangle=\delta_{n,j}$.
To show that the image of $-\tilde{\Delta}_{r}-\lambda I$ is closed, consider a sequence $g_{n}:=(-\tilde{\Delta}_{r}-\lambda I)(f_{n})$ such that $g_{n}\rightarrow g$. Similarly to before, we can conclude that since $g_{n}$ converges, then $f_{n}$ converges to some $f$, and thus $g=(-\tilde{\Delta}_{r}-\lambda I)(f)$. ∎
Corollary 3.8.
Let $(X,\mathsf{d},\mu)$ be a metric measure space such that for some $r_{0}>0$ both ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) hold for any $r\in(0,r_{0})$. Then
$\lim\limits_{r\downarrow 0}\big{(}\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\big{)}=+\infty.$
Proof.
Proposition 3.7 implies that
$\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}\leq\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r}).$ (41)
But for any $r>0$,
$\tilde{A}_{r}1=\frac{1}{2}\left(A_{r}1+A^{*}_{r}1\right)\geq\frac{1}{2}A_{r}1=\frac{1}{2}$
hence (41) implies that
$\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\geq\frac{1}{2r^{2}}\stackrel{{\scriptstyle r\downarrow 0}}{{\to}}+\infty.$
Let us provide our second spectral result on $\tilde{\Delta}_{r}$.
Proposition 3.9.
Let $(X,\mathsf{d},\mu)$ be a connected metric measure space satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$. Then the kernel of $\tilde{\Delta}_{r}$ contains constant functions only, and $\tilde{E}_{r}$ defines a scalar product on
$\mathrm{\Pi}(X,\mu)\coloneqq\left\{f\in L^{2}(X,\mu):\int_{X}f\mathop{}\!\mathrm{d}\mu=0\right\}.$ (42)
Proof.
Consider $f\in L^{2}(X,\mu)\backslash\{0\}$ such that $\tilde{\Delta}_{r}f=0$. Then $\tilde{E}_{r}(f)=0$. This implies that for $\mu$-a.e. $x\in X$,
[…]
which implies, in turn,
$\mu\left(\{y\in B_{r}(x):f(y)=f(x)\}\right)=\mu(B_{r}(x)).$ (44)
[…]
$A:=\{x\in F:\mu\left(\{y\in B_{r}(x):f(y)=f(x)\}\right)=\mu(B_{r}(x))\}.$ (45)
Then
$\mu(X\backslash A)=0,$ (46)
Since the measure of any open ball is positive, $X\backslash A$ cannot contain such a ball. This implies that $A$ is a dense subset of $X$. Take $z\in A$ and let $c=f(z)$. Consider
$I=\{x\in A:f(x)=c\},\quad\quad I^{\prime}=\{x\in A:f(x)\neq c\},$ (47)
and notice that $I\cup I^{\prime}=A$. Suppose by contradiction that $I^{\prime}\neq\emptyset$. Set
$W:=\bigcup_{x\in I}B_{r}(x),\quad\quad V:=\bigcup_{y\in I^{\prime}}B_{r}(y).$ (48)
Since $V$ and $W$ are open sets whose union contains $A$ which is dense in $X$, we must have
$W\cup V=X.$ (49)
Since $W$ and $V$ form an open cover of $X$, and $X$ is connected, if both $V$ and $W$ are different from the empty set, then there exist $x\in I$ and $y\in I^{\prime}$ such that
$B_{r}(x)\cap B_{r}(y)\neq\emptyset.$ (50)
However
$f|_{B_{r}(x)\cap B_{r}(y)}(w)$ $=f(x)\quad\quad\mu\text{-a.e. }w\in X,$ (51)
$f|_{B_{r}(x)\cap B_{r}(y)}(w)$ $=f(y)\quad\quad\mu\text{-a.e. }w\in X.$ (52)
This is not possible since $f(x)\neq f(y)$ and $\mu(B_{r}(x)\cap B_{r}(y))>0$. This implies that $I=A$ and that $f(w)=c$ for $\mu$-a.e. $w\in X$. Then $-\tilde{\Delta}_{r}$ has a non-trivial kernel consisting of the constant functions only. Moreover, since $-\tilde{\Delta}_{r}$ is non-negative (Remark 3.5), we get the desired property on (42).
We are now in a position to prove Theorem 1. We recall that the context of this statement is a compact uniformly locally doubling metric measure space $(X,\mathsf{d},\mu)$.
Proof.
For $k\geq 1$ integer, let $x_{0},\ldots,x_{k}\in X$ be distinct points. Set
$\bar{r}_{k}\coloneqq\min_{0\leq i\neq j\leq k}\frac{\mathsf{d}(x_{i},x_{j})}{4}\,\cdot$
For any $i\in\{0,\ldots,k\}$ and $y\in X$, define
$\tilde{f}_{i}(y)\coloneqq\left(1-\frac{\mathsf{d}(x_{i},y)}{\bar{r}_{k}}\right)^{+}\qquad\text{and}\qquad f_{i}(y)\coloneqq\frac{\tilde{f}_{i}(y)}{\|\tilde{f}_{i}\|_{2}}\,\cdot$
Note that each $f_{i}$ is an $L^{2}$-normalized Lipschitz function supported in $B_{\bar{r}_{k}}(x_{i})$, and that $(f_{0},\ldots,f_{k})$ is an orthonormal family of $L^{2}(X,\mu)$. Set
$V\coloneqq\mathrm{Span}(f_{0},\ldots,f_{k})\in\mathcal{G}_{k+1}(L^{2}(X,\mu))$
and observe that for any $r>0$,
$\tilde{\lambda}_{k,r}\leq\max_{\begin{subarray}{c}f\in V\\ \|f\|_{2}=1\end{subarray}}\tilde{E}_{r}(f).$
Consider $r\in(0,\bar{r}_{k})$ and $i\neq j$ in $\{0,\ldots,k\}$. If $x\in B_{2\bar{r}_{k}}(x_{i})$, then $f_{j}(y)=0$ for any $y\in B_{r}(x)$, while if $x\notin B_{2\bar{r}_{k}}(x_{i})$, then $f_{i}(y)=0$ for any $y\in B_{r}(x)$. In both cases,
$(f_{i}(x)-f_{i}(y))(f_{j}(x)-f_{j}(y))=0$
for any $y\in B_{r}(x)$. Thus
$\tilde{E}_{r}(f_{i},f_{j})=0.$ (53)
Consider $f\in V$ such that $\|f\|_{2}=1$. Then $f=\sum_{i=0}^{k}a_{i}f_{i}$ for some $a_{0},\ldots,a_{k}\in\mathbb{R}$ such that $\sum_{i=0}^{k}a_{i}^{2}=1$. By (53), we get
$\tilde{E}_{r}(f)=\sum_{i=0}^{k}a_{i}^{2}\tilde{E}_{r}(f_{i})\leq\max_{0\leq i\leq k}\tilde{E}_{r}(f_{i}).$
Let $r_{0},C$ be the parameters of the local doubling property of $(X,\mathsf{d},\mu)$, see Definition 2.1. Then for any $i\in\{0,\ldots,k\}$ and $r\in(0,r_{0})$,
$\tilde{E}_{r}(f_{i})$ $\leq\frac{\operatorname{Lip}^{2}(f_{i})}{4}\int_{X}\fint_{B_{r}(x)}\underbrace{\left(1+\frac{V(x,r)}{V(y,r)}\right)}_{\leq 1+C^{2}}\underbrace{\frac{\mathsf{d}^{2}(x,y)}{r^{2}}}_{\leq 1}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$\leq\frac{\operatorname{Lip}^{{2}}(f_{i})(1+C^{2})\mu(X)}{4}\,\cdot$
Therefore, for any $r<\tilde{r}_{k}\coloneqq\min(\bar{r}_{k},R)$, we get
$\tilde{\lambda}_{k,r}\leq\tilde{C}\coloneqq\frac{(1+C^{2})\mu(X)}{4}\max_{0\leq i\leq k}\operatorname{Lip}^{{2}}(f_{i})\,\cdot$ (54)
By Corollary 3.8, there exists $r_{k}\in(0,\tilde{r}_{k})$ such that $\tilde{C}<\min\sigma_{\mathrm{ess}}(-\tilde{\Delta}_{r})$ for any $r\in(0,r_{k})$. Then Remark 3.6 implies that for such an $r$ the operator $-\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues $\lambda_{0}(-\tilde{\Delta}_{r})\leq\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$ such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$. That $\lambda_{0}(-\tilde{\Delta}_{r})=0$ follows from Proposition 3.9. Moreover, by Remark 3.6 and Proposition 3.9, we know that
$\tilde{\lambda}_{1,r}=\min_{f\in\Pi(X,\mu)}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}}$
where $\Pi(X,\mu)$ is as in (42). Since $-\Delta_{r}$ has a kernel which is $L^{2}$-orthogonal to $\Pi(X,\mu)$ (Proposition 3.9), we have $\tilde{E}_{r}(f)>0$ for any $f\in\Pi(X,\mu)$, hence we get
$\tilde{\lambda}_{1,r}>0.$
4. First eigenvalue of torus and hypercubes
In this section, we derive some results which will be applied in Section 6. We let $m$ be a positive integer kept fixed throughout the section.
4.1. A preliminary lemma
We begin with a result where we use the normalized sinc function, namely
$\operatorname{sinc}(\rho)\coloneqq\begin{cases}\displaystyle\frac{\sin(\pi\rho)}{\pi\rho}\qquad&\rho\in\mathbb{R}\backslash\{0\},\\ 1&\rho=0,\end{cases}$
and the following notation: for any $p=(p_{1},...,p_{m})\in\mathbb{Z}^{m}$,
$J(p)\coloneqq\{i\in\{1,\ldots,m\}:p_{i}\neq 0\},\qquad j(p)\coloneqq\#J(p).$ (55)
Lemma 4.1.
$\liminf_{r\rightarrow 0}\inf_{0\neq p\in\mathbb{Z}^{m}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$ (56)
Proof.
We start by pointing out that for any $r>0$ and $p\in\mathbb{Z}^{m}\backslash\{0\}$,
$\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\neq 1,$
so that
$\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$ (57)
Moreover, for any $r>0$, if $|p|_{\infty}\coloneqq\max_{1\leq i\leq m}|p_{i}|\to+\infty$, then
$\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|\rightarrow\frac{1}{r^{2}}>0,$
hence there exists $R>0$ such that
$\inf_{0\neq p\in\mathbb{Z}^{m}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|=\min_{\begin{subarray}{c}0\neq p\in\mathbb{Z}^{m}\\ |p|_{\infty}<R\end{subarray}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$
If (56) were to fail, due to the previous line, there would exist sequences $(r_{n})\subset(0,+\infty)$ and $(p^{(n)})\subset\mathbb{Z}^{m}\backslash\{0\}$ such that $r_{n}\to 0$ and
$\lim_{n}\left|\frac{1}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p^{(n)}_{i}r)\right)\right|=0.$ (58)
We have two cases.
  • There exists $\alpha>0$ and $j\in\{1,\ldots,m\}$ such that $\liminf_{n}p^{(n)}_{j}r_{n}>\alpha$.
  • For any $j\in\{1,\ldots,m\}$, one has $\liminf_{n}p^{(n)}_{j}r_{n}=0$.
If the first one were true, then we would have
$\liminf_{n}\left|1-\prod_{i\in J(p)}\operatorname{sinc}(p^{(n)}_{i}r)\right|>|1-\alpha|,$ (59)
and so (58) couldn’t hold. On the contrary, if the second case were true, up to extracting a subsequence we would have $p^{(n)}_{j}r_{n}\to 0$ for any $j\in\{1,\ldots,m\}$. For any $y=(y_{1},\ldots,y_{m})\in\mathbb{R}^{m}$, set
$G(y)\coloneqq 1-\prod_{j=1}^{m}\operatorname{sinc}(y_{j}).$ (60)
Then $G$ is smooth on $\mathbb{R}^{m}$ and satisfies
$G(y)=\frac{1}{2}|y|^{2}+o(|y|^{3}),\qquad|y|\to 0.$ (61)
As a consequence, for $y\in\mathbb{R}^{m}$ such that $|y|$ is small enough,
$\left|G(y)\right|\geq\frac{1}{4}|y|^{2}.$ (62)
Then, for large enough $n$,
$\left|1-\prod_{i\in I(p^{(n)})}\operatorname{sinc}(p^{(n)}_{i}r_{n})\right|$ $=G(r_{n}p^{(n)})\geq\frac{1}{4}|r_{n}p^{(n)}|^{2}\geq\frac{1}{4}r_{n}^{2},$
because $p^{(n)}\neq 0$ implies that there exists at least one $i$ such that $|p^{(n)}_{i}|\geq 1$. Thus we obtain
$\lim_{n}\left|\frac{1}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(\pi p^{(n)}_{i}r_{n})\right)\right|\geq\frac{1}{4}$ (63)
and so (58) could not hold. This concludes the proof. ∎
4.2. Torus
Consider the torus
$\mathbb{T}^{m}:=\mathbb{R}^{m}/\big{(}-1+2\mathbb{Z}\big{)}^{m}$
with its natural quotient map $\pi:\mathbb{R}^{m}\to\mathbb{T}^{m}$. Let $\pi^{-1}$ be the inverse of the bijective map
$\pi:[-1,1)^{m}\to\mathbb{T}^{m}.$
Let $\mathsf{d}_{\infty}$ be the distance in $\mathbb{R}^{m}$ associated with the infinity norm, given by
$\mathsf{d}_{\infty}(x,y):=\max\{|x_{i}-y_{i}|:i\in\{1,...,m\}\}$ (64)
for any $x=(x_{1},\ldots,x_{m})$ and $y=(y_{1},\ldots,y_{m})$ in $\mathbb{R}^{m}$. With respect to this distance, the open ball of radius $r>0$ centered at $x\in\mathbb{R}^{m}$ is
$Q_{r}(x):=\prod_{i=1}^{m}(-r+x_{i},x_{i}+r).$ (65)
For any $x,y\in\mathbb{T}^{m}$, set
$\tilde{\mathsf{d}}_{\infty}(x,y)=\inf_{z\in\pi^{-1}(x),w\in\pi^{-1}(y)}\mathsf{d}_{\infty}(z,w).$
Then $\tilde{\mathsf{d}}_{\infty}$ defines a distance on $\mathbb{T}^{m}$, and we denote by $\tilde{Q}_{r}(x)$ the open ball of radius $r>0$ centered at $x\in\mathbb{T}^{m}$ with respect to this distance. We also introduce the probability measure
$\mathbb{L}^{m}\coloneqq\pi_{\#}\left(\frac{\mathcal{L}^{m}}{2^{m}}\right)$
on $\mathbb{T}^{m}$. Note that this is also the normalized Haar measure of $\mathbb{T}^{m}$ seen as a Lie group.
It is obvious that the metric measure space $\mathfrak{T}^{m}\coloneqq(\mathbb{T}^{m},\tilde{\mathsf{d}}_{\infty},\mathbb{L}^{m})$ satisfies the assumptions of Theorem 1, hence we know that for any small enough $r$,
$\tilde{\lambda}_{1,r}=\lambda_{1}(-\tilde{\Delta}_{r})>0.$
Then the following holds.
Proposition 4.2.
$\liminf_{r\rightarrow 0}\lambda_{1}(-\tilde{\Delta}_{r})>0.$ (66)
Proof.
For any $x\in\mathbb{T}^{m}$ and $r>0$ small enough, the ball $\tilde{Q}_{r}(x)\subset\mathbb{T}^{m}$ is given by
$\tilde{Q}_{r}(x)=\pi(Q_{r}(\tilde{x}))$ (67)
where $\tilde{x}$ is any element in $\pi^{-1}(x)$, and $Q_{r}(\tilde{x})$ is as in (65). Then
$\mathbb{L}^{m}(\tilde{Q}_{r}(x))=\frac{\mathcal{L}^{m}(Q_{r}(\tilde{x}))}{2^{m}}=r^{m}$ (68)
so that the $r$-energy functional of $\mathfrak{T}^{m}$ writes as
$\tilde{E}_{r,\mathfrak{T}^{m}}(f)=\frac{1}{r^{m}}\int_{\mathbb{T}^{m}}\int_{\tilde{Q}_{r}(x)}\left(\frac{f(x)-f(y)}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$ (69)
for any $f\in L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$.
We act by contradiction. Assume that there exist $r_{n}\rightarrow 0$ and $\{f_{n}\}\subset L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$ satisfying $\int_{\mathbb{T}^{m}}f_{n}\mathop{}\!\mathrm{d}\mathbb{L}^{m}=0$ and $\|f_{n}\|_{L^{2}(\mathbb{T}^{m})}=1$, such that
$-\tilde{\Delta}_{r_{n}}f_{n}=\lambda_{1}(-\tilde{\Delta}_{r_{n}})f_{n},\quad\quad\lambda_{1}(-\tilde{\Delta}_{r_{n}})\rightarrow 0.$ (70)
With no loss of generality, we assume that each $r_{n}$ is small enough to ensure that $\tilde{E}_{r_{n},\mathfrak{T}^{m}}$ writes as in (69).
From (68) we can write, for any $n$ and $\mathbb{L}^{m}$-a.e. $x\in\mathbb{T}^{m}$,
$-\tilde{\Delta}_{r_{n}}f_{n}(x)$ $=\frac{1}{r_{n}^{2}}\left(f_{n}(x)-\frac{1}{r_{n}^{m}}\int_{\mathbb{T}^{m}}1_{\tilde{Q}_{r_{n}}(x)}(y)f_{n}(y)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)\right)$
$=\frac{1}{r_{n}^{2}}\left(f_{n}(x)-\frac{1}{r_{n}^{m}}\bigg{(}1_{\tilde{Q}_{r_{n}}(0)}*f_{n}\bigg{)}(x)\right).$ (71)
We consider the Fourier decomposition of $f_{n}$, $1_{\tilde{Q}_{r_{n}}(0)}$, and $-\tilde{\Delta}_{r_{n}}f_{n}$, namely
$f_{n}=\sum_{p\in\mathbb{Z}^{m}}a_{p,n}e_{p}$
$1_{\tilde{Q}_{r_{n}}(0)}=\sum_{p\in\mathbb{Z}^{m}}b_{p,n}e_{p}$
$-\tilde{\Delta}_{r_{n}}f_{n}=\sum_{p\in\mathbb{Z}^{m}}c_{p,n}e_{p}$
where $\{e_{p}\}_{p\in\mathbb{Z}^{m}}$ is the orthonormal basis of $L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$ given by
$e_{p}:\mathbb{T}^{m}\ni x\mapsto e^{i\pi p\cdot\pi^{-1}(x)}\qquad\forall\,p\in\mathbb{Z}^{m}.$
Since the Fourier coefficients of a convolution are the product of the coefficients, we obtain from (71) that
$c_{p,n}=\frac{1}{r_{n}^{2}}\left(a_{p,n}-\frac{b_{p,n}}{r_{n}^{m}}a_{p,n}\right)=\frac{a_{p,n}}{r_{n}^{2}}\left(1-\frac{b_{k,n}}{r_{n}^{m}}\right).$ (72)
We can compute each coefficient $b_{p,n}$ by means of Fubini’s Theorem; we obtain
$b_{p,n}=\int_{Q_{r_{n}}(0)}e^{i\pi p\cdot x}\frac{d\mathcal{L}^{m}(x)}{2^{m}}=r_{n}^{m}\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n}).$ (73)
Thus
$c_{p,n}=\frac{a_{p,n}}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n})\right).$ (74)
Using Lemma 4.1, for $\0p\in\mathbb{Z}\backslash\{0\}$ we conclude that there exists $\alpha>0$ such that
$\frac{1}{r_{n}^{2}}\left|1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n})\right|\geq\alpha.$ (75)
This implies that
$|c_{p,n}|^{2}\geq|a_{p,n}|^{2}\alpha^{2}.$ (76)
By Parseval’s identity, and since $f_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$ we have $a_{0,n}=0$,
$\|-\tilde{\Delta}_{r_{n}}f_{n}\|_{L^{2}(\mathbb{T}^{m})}^{2}=\sum_{p\in\mathbb{Z}^{m}}|c_{p,n}|^{2}\geq\sum_{p\in\mathbb{Z}^{m}\backslash\{0\}}|a_{p,n}|^{2}\alpha^{2}=\alpha^{2}\|f_{n}\|_{L^{2}(\mathbb{T}^{m})}^{2},$ (77)
in contradiction with $\|-\tilde{\Delta}_{r_{n}}f_{n}\|_{L^{2}(\mathbb{T}^{m})}\rightarrow 0$ provided by (70). ∎
4.3. Shrinking hypercubes
For any $b>0$, consider the metric measure space $\mathfrak{Q}^{m}(b)\coloneqq([0,b]^{m},\mathsf{d}_{\infty},\mathcal{L}^{m})$. It trivially satisfies the assumptions of Theorem 1, so that for any small enough $r$,
$\tilde{\lambda}_{1,r}=\lambda_{1}(-\tilde{\Delta}_{r,\mathfrak{Q}^{m}(b)})>0.$ (78)
Then the following holds.
Lemma 4.3.
$\liminf_{b\rightarrow 0}\lim_{r\to 0}\lambda_{1}(-\tilde{\Delta}_{r,\mathfrak{Q}^{m}(b)})=+\infty.$ (79)
Proof.
We suppose by contradiction that (79) fails. Then there exist $b_{n}\rightarrow 0$ and $r_{n}\rightarrow 0$ such that
$\lim_{n}\lambda_{1}(-\tilde{\Delta}_{r_{n},\mathfrak{Q}^{m}(b)})<+\infty.$
Since we are first taking the limit in $r$ and then in $b$, we can assume that $\overline{r}_{n}:=\frac{r_{n}}{b_{n}}\rightarrow 0$. For any $n$, by a simple scaling argument we have
$\lambda_{1}(-\tilde{\Delta}_{r_{n},\mathfrak{Q}^{m}(b_{n})})=\frac{1}{b_{n}^{2}}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})$
thus
$\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\rightarrow 0.$ (80)
From (78), assuming that each $r_{n}$ is small enough, we know that there exists $f_{n}\in\Pi([0,1]^{m},\mathcal{L}^{m})$ such that $\|f_{n}\|_{L^{2}([0,1]^{m},\mathcal{L}^{m})}=1$ and
$\tilde{E}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}(f_{n})=\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}).$
Consider the continuous function
$\begin{array}[]{cccccc}T&:&\mathbb{T}^{m}&\rightarrow&[0,1]^{m}\\ &&x&\mapsto&(|\bar{x}_{1}|,...,|\bar{x}_{m}|)\end{array}$
where $\bar{x}=\pi^{-1}(x)\in[-1,1)^{m}$. For any $n$, set
$\tilde{f}_{n}=f_{n}\circ T\in L^{2}(\mathbb{T}^{m},\mathbb{L}^{m}).$
From this we have that $f_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$ Let us prove that
$\|\tilde{f}_{n}\|_{L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})}=1.$
Let $C_{1},\ldots,C_{2^{m}}$ denote the $2^{m}$ sets of the form $I_{1}\times\ldots\times I_{m}$ where each $I_{i}$ is either $[-1,0]$ or $[0,1]$. For any $j\in\{1,\ldots,2^{m}\}$, set
$\begin{array}[]{cccccc}N_{j}&:&[0,1]^{m}&\rightarrow&C_{j}\\ &&(\xi_{1},\ldots,\xi_{m})&\mapsto&(\varepsilon_{i}\xi_{1},\ldots,\varepsilon_{m}\xi_{m})\end{array}$
where $\varepsilon_{i}$ is $1$ if $I_{i}=[0,1]$ and $-1$ otherwise. Note that $N_{j}$ is an isometry which preserves the Lebesgue measure, and that $T\circ\pi\circ N_{j}$ is equal to the identity. Then
$\|\tilde{f}_{n}\|_{L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})}^{2}=\int_{\mathbb{T}^{m}}(f_{n}\circ T)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}$ $=\frac{1}{2^{m}}\int_{[-1,1]^{m}}(f_{n}\circ T\circ\pi)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
$=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{C_{j}}(f_{n}\circ T\circ\pi)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
$=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{[0,1]^{m}}(f_{n}\circ T\circ\pi\circ N_{j})^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
$=\|f_{n}\|_{L^{2}([0,1]^{m},\mathcal{L}^{m})}^{2}=1.$ (81)
We claim that
$\tilde{E}_{\overline{r}_{n},\mathfrak{T}^{m}}(\tilde{f}_{n})\leq 2^{m}\tilde{E}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}(f_{n}).$ (82)
Since $\tilde{f}_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$, the latter provides a contradiction with Proposition 4.2, namely
$0<\lambda_{1}(-\tilde{\Delta}_{\bar{r}_{n},\mathfrak{T}^{m}})\leq 2^{m}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\to 0$
Given $x\in\mathbb{T}^{m}$, define
$\tilde{G}_{n}(x)\coloneqq\frac{1}{\overline{r}_{n}^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{\tilde{f}_{n}(x)-\tilde{f}_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y).$
For any $x\in[0,1]^{m}$, set
${G}_{n}(x)$ $\coloneqq\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{1}{V(x,\overline{r}_{n})}+\frac{1}{V(y,\overline{r}_{n})}\right)$
×$\qquad\qquad\quad\times\left(\frac{f_{n}(x)-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$
where
$V(z,r)\coloneqq\mathcal{L}^{m}(Q_{r}(z)\cap[0,1]^{m})$ (83)
for any $z\in[0,1]^{m}$. Then
$4\tilde{E}_{\overline{r}_{n},\mathfrak{T}^{m}}(\tilde{f}_{n})$ $=\int_{\mathbb{T}^{m}}\tilde{G}_{n}(x)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$
$4\tilde{E}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}(f_{n})$ $=\int_{[0,1]^{m}}{G}_{n}(x)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x).$
Moreover, for all $x\in\mathbb{T}^{m}$,
$\frac{1}{\overline{r}_{n}^{m}}\leq\frac{2^{m}}{\mathcal{L}^{m}(Q_{\overline{r}_{n}}(T(x))\cap[0,1]^{m})}=\frac{2^{m}}{V(T(x),\overline{r}_{n})}.$ (84)
For any $x\in\mathbb{T}^{m}$, there exists some $k\in\{1,...,2^{m}\}$ and $\overline{x}\in C_{k}$ such that $\pi(\overline{x})=x$. We will now consider for each $j\in\{1,...,2^{m}\}$ the rectangle given by
$R_{n,j}(x):=T(\pi(C_{j})\cap\tilde{Q}_{\overline{r}_{n}}(x))\subset[0,1]^{m}.$ (85)
and we point out that
$R_{n,j}(x)\subset R_{n,k}(x)=[0,1]^{m}\cap Q_{\overline{r}_{n}}(T(x))$ (86)
for any $j\in\{1,...,2^{m}\}$. This follows since for each $i\in\{1,...,m\}$ we have 4 possibilities
  1. (1)
    $|\overline{x}_{i}|<\overline{r}_{n}$
  2. (2)
    $|1-\overline{x}_{i}|<\overline{r}_{n}$
  3. (3)
    $|-1-\overline{x}_{i}|<\overline{r}_{n}$
  4. (4)
    $\neg\left((1)\lor(2)\lor(3)\right)$.
If $\pi_{i}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the projection in the $i$-th coordinate, we conclude
$\pi_{i}(R_{n,k}(x))=\begin{cases}[0,\overline{r}_{n}+|\overline{x}_{i}|]&\text{ if }(1)\\ [-\overline{r}_{n}+|\overline{x}_{i}|,1]&\text{ if }(2)\\ [-\overline{r}_{n}+|\overline{x}_{i}|,1]&\text{ if }(3)\\ [-\overline{r}_{n},\overline{r}_{n}]&\text{ if }(4)\end{cases}$ (87)
$\pi_{l}(R_{n,k}(x))=\begin{cases}\pi_{i}(R_{n,l}(x))=\pi_{i}(R_{n,k}(x))&\text{ if }\pi_{i}(C_{j})=\pi_{i}(C_{k})\\ [0,\overline{r}_{n}-|\overline{x}_{i}|]&\text{ if }(1)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ [2-\overline{r}_{n}-|\overline{x}_{i}|,1]&\text{ if }(2)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ [2-\overline{r}_{n}-|\overline{x}_{i}|,1]&\text{ if }(3)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ \emptyset&\text{ if }(4)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right).\end{cases}$ (88)
Thus we conclude that for all $i\in\{1,...,m\}$ and $l\in\{1,...,2^{m}\}$ we have $\pi_{i}(R_{n,l}(x))\subset\pi_{i}(R_{n,k}(x))$, and since these sets are rectangles, we conclude equation (86).
From (84) we can deduce
$\tilde{G}_{n}(x)$ $=\frac{1}{\bar{r}_{n}^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{\tilde{f}_{n}(x)-\tilde{f}_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (89)
$\leq\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)$ (90)
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (91)
$=\sum_{j=1}^{2^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)\cap\pi(C_{j})}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)$ (92)
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (93)
$=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j}}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(\pi(y)),\overline{r}_{n})}\right)$ (94)
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(y)))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (95)
For each integral, change coordinates by $N_{j}$ to conclude
$\tilde{G}_{n}(x)$ $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{N_{j}^{-1}(\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j})}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(\pi(N_{j}(y))),\overline{r}_{n})}\right)\times$ (96)
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(N_{j}(y))))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y).$ (97)
We have that
$N_{j}^{-1}(\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j})=R_{n,j}(x),$ (98)
so by equation (86) and the fact that $T\circ\pi\circ N_{j}=id$, we conclude
$\tilde{G}_{n}(x)$ $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{R_{n,j}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)\times$ (99)
×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (100)
$\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{R_{n,k}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(y,\overline{r}_{n})}\right)\times$ (101)
×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (102)
$=\int_{[0,1]^{m}\cap Q_{\overline{r}_{n}}(T(x))}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(y,\overline{r}_{n})}\right)\times$ (103)
×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (104)
$=2^{m}G_{n}(T(x)).$ (105)
Now we integrate both sides in $\mathbb{T}^{m}$ and change variables by $\pi$ and $N_{j}$ to conclude
$4\tilde{E}_{\overline{r}_{n},\mathfrak{T}^{m}}(\tilde{f}_{n})=\int_{\mathbb{T}^{m}}\tilde{G}_{n}(x)\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$ $\leq 2^{m}\int_{\mathbb{T}^{m}}G_{n}(T(x))\mathop{}\!\mathrm{d}\mathbb{L}^{m}(x)$ (106)
$=\int_{[-1,1)^{m}}G_{n}(T(\pi(x)))\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (107)
$=\sum_{j=1}^{2^{m}}\int_{C_{j}}G_{n}(T(\pi(x)))\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (108)
$=\sum_{j=1}^{2^{m}}\int_{[0,1]^{m}}G_{n}(T(\pi(N_{j}(x))))\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (109)
$=\sum_{j=1}^{2^{m}}\int_{[0,1]^{m}}G_{n}(x)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (110)
$=2^{m}4\tilde{E}_{\overline{r}_{n},[0,1]^{m}}(f_{n})=2^{m}4\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)}).$ (111)
With this we obtain (82). ∎
5. $L^{2}$ convergence
In this section, we prove Theorem 2. Let $M$ be a smooth, compact, connected manifold of dimension $m\geq 2$. Assume that $M$ is endowed with a smooth Riemannian metric $g$ and let $\mathsf{d}_{g}$ and $\mathrm{vol}_{g}$ be the associated Riemannian distance and volume measure on $M$.
In this smooth context, the function $V(\cdot,r)$ is obviously continuous for any $r>0$. Since $M$ is compact, this implies that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$). Then Lemma 3.2 applies and ensures that $\tilde{\Delta}_{r}$ is a bounded self-adjoint operator acting on $L^{2}(M,\mathrm{vol}_{g})$. The compactness of $M$ also ensures that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is locally Ahlfors regular: there exists a constant $C>1$ such that for any $x\in M$ and $r\in(0,\mathrm{diam}(M)]$,
$C^{-1}r^{m}\leq V(x,r)\leq Cr^{m}.$ (112)
Note that this condition trivially implies a uniform local doubling property for $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$.
5.1. Convergence in the sense of distributions
Recall that $C_{m}$ is defined in (3). For any $x\in M\backslash\partial M$, we let $\exp_{x}$ be the exponential map centered at $x$. We identify $T_{x}M$ with $\mathbb{R}^{m}$ and write $\mathbb{B}^{m}_{r}(v)$ for the Euclidean ball in $\mathbb{R}^{m}$ centered at $v$ with radius $r>0$. Then there exists $\delta>0$ such that the restriction of $\exp_{x}$ to $\mathbb{B}^{m}_{\delta}(0)$ is a diffeomorphism onto its image; recall that the injectivity radius $\mathrm{i}_{M}(x)$ of $M$ at $x$ is the supremum of the set of such numbers $\delta$. We let $J_{x}$ be the Radon-Nikodym derivative of the measure $(\exp_{x}^{-1})_{\#}\mathrm{vol}_{g}$ with respect to the Lebesgue measure $\mathcal{L}^{m}$. It is well-known that for any $\xi\in\mathbb{B}_{\mathrm{i}_{M}}^{m}(0)$,
$J_{x}(\xi)=1+\underline{O}_{K}(|\xi|^{2})$
where for any $h>0$, the notation $\underline{O}_{K}(h)$ stands for a quantity independent on $x\in K$ whose absolute value divided by $h$ is bounded. Here $K$ is a compact subset of $M$. We write $\underline{O}$ instead of $\underline{O}_{M}$. Then the following holds.
Proposition 5.1.
Consider $f,\psi\in C^{2}(M)$. Then
$\lim_{r\rightarrow 0}\langle-\tilde{\Delta}_{r}f,\psi\rangle_{2}=C_{m}\int_{M}\langle\mathop{}\!\mathrm{d}f,\mathop{}\!\mathrm{d}\psi\rangle_{g}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}.$ (113)
Proof.
For any $x\in M$ and $r>0$, set
$\tilde{e}_{r}(f,\psi;x)\coloneqq\frac{1}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
so that
$\langle-\tilde{\Delta}_{r}f,\psi\rangle_{2}=\int_{M}\tilde{e}_{r}(f,\psi;y)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y).$
On one hand,
$|\tilde{e}_{r}(f,\psi;x)|$ $\leq\frac{1}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\frac{|f(x)-f(y)||\psi(x)-\psi(y)|}{r^{2}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$\leq\frac{1}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\frac{\operatorname{Lip}(f)\operatorname{Lip}(\psi)\mathsf{d}_{g}^{2}(x,y)}{r^{2}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$\leq\frac{\operatorname{Lip}(f)\operatorname{Lip}(\psi)}{4}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y).$
By (112), we obtain
$|\tilde{e}_{r}(f,\psi;x)|\leq\frac{\operatorname{Lip}(f)\operatorname{Lip}(\psi)(1+C^{2})}{4}\,\cdot$ (114)
On the other hand, assume that $r$ is smaller than $\mathrm{i}_{M}(x)$, and consider $\tilde{f}\coloneqq f\circ\exp_{x}$ and $\tilde{\psi}\coloneqq\psi\circ\exp_{x}$ on $\mathbb{B}^{m}_{r}(0)$. The first-order Taylor expansion of $\tilde{f}$ and $\tilde{\psi}$ yields
$\tilde{f}(\xi)$ $=\tilde{f}(0)+(\mathop{}\!\mathrm{d}\tilde{f})_{0}(\xi)+\underline{O}_{\{0\}}(|\xi|^{2})$
$\tilde{\psi}(\xi)$ $=\tilde{\psi}(0)+(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}(\xi)+\underline{O}_{\{0\}}(|\xi|^{2}).$
Then
$\phantom{=}\int_{B_{r}(x)}(f(x)-f(y))(\psi(x)-\psi(y))\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$=\int_{\mathbb{B}^{m}_{r}(0)}(\tilde{f}(0)-\tilde{f}(\xi))(\tilde{\psi}(0)-\tilde{\psi}(\xi))J(\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$=\int_{\mathbb{B}^{m}_{r}(0)}((\mathop{}\!\mathrm{d}\tilde{f})_{0}(\xi)+\underline{O}_{\{0\}}(r^{2}))((\mathop{}\!\mathrm{d}\tilde{\psi})_{0}(\xi)+\underline{O}_{\{0\}}(r^{2}))(1+\underline{O}_{\{0\}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$=\sum_{i,j=1}^{m}[(\mathop{}\!\mathrm{d}\tilde{f})_{0}]_{i}[(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}]_{j}\int_{\mathbb{B}^{m}_{r}(0)}\xi_{j}\xi_{i}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{\{0\}}(r^{m+3})$
$=(\mathop{}\!\mathrm{d}\tilde{f})_{0}\cdot(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}\int_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{\{0\}}(r^{m+3}).$
Moreover, it is known that (see e.g. [MT23, Remark 2.11])
$1+\frac{V(x,r)}{V(y,r)}={2}+\underline{O}_{\{x\}}(r^{2})$
and since $V(x,r)/\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0)))\to 1$ as $r\downarrow 0$, we obtain
$\phantom{=}\fint_{B_{r}(x)}\left(1+\frac{V(x,r)}{V(y,r)}\right)(f(x)-f(y))(\psi(x)-\psi(y))\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$
$=\frac{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0))({2}+\underline{O}_{\{x\}}(r^{2}))}{V(x,r)}\fint_{\mathbb{B}^{m}_{r}(0)}(\tilde{f}(0)-\tilde{f}(\xi))(\tilde{\psi}(0)-\tilde{\psi}(\xi))J(\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$=(2+\underline{O}_{\{x\}}(r^{2}))\left((\mathop{}\!\mathrm{d}\tilde{f})_{0}\cdot(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}\fint_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{\{0\}}(r^{3})\right).$
Since $(\mathop{}\!\mathrm{d}\tilde{f})_{0}\cdot(\mathop{}\!\mathrm{d}\tilde{\psi})_{0}=\langle\mathop{}\!\mathrm{d}f,\mathop{}\!\mathrm{d}\psi\rangle_{g}(x)$ and
$\fint_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)=2r^{2}C_{m}$
by change of variable $\xi\leftrightarrow\eta/r^{2}$, we eventually obtain that
$\tilde{e}_{r}(f,\psi;x)=C_{m}\langle\mathop{}\!\mathrm{d}f,\mathop{}\!\mathrm{d}\psi\rangle_{g}(x)+\underline{O}_{\{x\}}(r)\qquad\text{as $r\to 0$.}$ (115)
By (114) and the compactness of $M$, we can apply the dominated convergence theorem to the functions $\tilde{e}_{r}(f,\psi;\cdot)$. Then we get (113). ∎
Using integration by parts in (113), we immediately obtain the following.
Corollary 5.2.
Let $\partial g$ be the Riemannian metric induced by $g$ on $\partial M$. For any $f\in\mathcal{C}^{\infty}(M)$, the following convergence holds in the sense of distributions as $r\downarrow 0$:
$(\tilde{\Delta}_{r}f)\,\mathrm{vol}_{g}\to C_{m}\bigg{(}(\Delta_{g}f)\,\mathrm{vol}_{g}+(\partial_{\nu}^{g}f)\,\mathrm{vol}_{\partial g}\bigg{)}.$
5.2. Pointwise convergence
We aim to prove Theorem 2 in a similar way as Proposition 5.1, that is to say, by means of the dominated convergence theorem. To this aim, we first establish that pointwise convergence holds $\mathrm{vol}_{g}$-a.e. on $M$. We recall that $\partial M$ is a $\mathrm{vol}_{g}$-negligible subset of $M$.
Proposition 5.3.
Let $f\in C^{\infty}(M)$. Then for any $x\in M-\partial M$,
$\lim_{r\rightarrow 0}\tilde{\Delta}_{r}f(x)=C_{m}\Delta_{g}f(x).$ (116)
Moreover, the convergence is uniform on any compact subset of $M-\partial M$.
Proof.
Let $K$ be a compact subset of $M-\partial M$. Consider $x\in K$ and $r\in(0,\mathrm{i}_{M}(x))$. Set $\tilde{f}_{x}\coloneqq f\circ\exp_{x}$. Acting like in the proof of Proposition 5.1, we get
$\tilde{\Delta}_{r}f(x)=\frac{(2+\underline{O}_{K}(r^{2}))}{2r^{2}}\fint_{\mathbb{B}^{m}_{r}(0)}\left(\tilde{f}_{x}(\xi)-\tilde{f}_{x}(0)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi).$ (117)
The second-order Taylor expansion of $\tilde{f}_{x}$ yields
$\tilde{f}_{x}(\xi)$ $=\tilde{f}_{x}(0)+(\mathop{}\!\mathrm{d}\tilde{f}_{x})_{0}(\xi)+\frac{1}{2}(\mathop{}\!\mathrm{d}^{(2)}\tilde{f}_{x})_{0}(\xi,\xi)+\underline{O}_{K}(|\xi|^{3})$
hence we get
$\fint_{\mathbb{B}^{m}_{r}(0)}\tilde{f}_{x}(\xi)-\tilde{f}_{x}(0)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$ $=\fint_{\mathbb{B}^{m}_{r}(0)}(\mathop{}\!\mathrm{d}\tilde{f}_{x})_{0}(\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)$
$+\frac{1}{2}\fint_{\mathbb{B}^{m}_{r}(0)}(\mathop{}\!\mathrm{d}^{(2)}\tilde{f}_{x})_{0}(\xi,\xi)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)+\underline{O}_{K}(r^{3}).$
The first term vanishes by symmetry. The second term is equal to
$\frac{1}{2}\Delta\tilde{f}_{x}(0)\fint_{\mathbb{B}^{m}_{r}(0)}\xi_{1}^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\xi)=\Delta_{g}f(x)r^{2}C_{m}.$
In the end we get
$\tilde{\Delta}_{r}f(x)$ $=\frac{2+\underline{O}_{K}(r^{2})}{2r^{2}}\left(\Delta_{g}f(x)r^{2}C_{m}+\underline{O}_{K}(r^{3})\right)$
$=(1+\underline{O}_{K}(r^{2}))\left(C_{m}\Delta_{g}f(x)+\underline{O}_{K}(r)\right)$
from which follows the desired result, by letting $r\downarrow 0$. ∎
5.3. Uniform bound
We wish now to provide a uniform $L^{\infty}$ bound for the functions $-\tilde{\Delta}_{r}\psi$ where $r$ is in a neighborhood of zero.
Let us first consider the case $\partial M=\emptyset$. From Proposition 5.3, we have uniform convergence
$\|\tilde{\Delta}_{r}f-C_{m}\Delta_{g}f\|_{\infty}\to 0$
so that
$\|\tilde{\Delta}_{r}f-C_{m}\Delta_{g}f\|_{2}\leq\|\tilde{\Delta}_{r}f-C_{m}\Delta_{g}f\|_{\infty}\mathrm{vol}_{g}(M)\to 0.$
Let us now deal with the case $\partial M\neq\emptyset$.
Proposition 5.4.
Assume that $\partial M\neq\emptyset$. Consider $f\in C^{\infty}_{\nu}(M)$. Then there exists $r_{0}>0$ such that
$\sup_{0<r<r_{0}}\|\tilde{\Delta}_{r}f\|_{L^{\infty}(M)}<+\infty.$ (118)
Proof.
Since $\partial M$ is compact, we can find a finite collection of smooth parameterizations $\psi_{i}:(-4,4)^{m-1}\rightarrow\partial M$ such that
$\partial M=\bigcup_{i}\psi_{i}\left(\left[-1,1\right]^{m-1}\right).$ (119)
Step 1.
We work with any of the previous $\psi_{i}$ which we denote by $\psi$. For any $x\in\partial M$, let $\nu_{x}\in T_{x}M$ be the unit inner normal vector of $\partial M$ at $x$. Since $\partial M$ is smooth, there exists $\epsilon>0$ such that the map $E:\partial M\times[0,\epsilon]\rightarrow M$ given by
$E(x,t)=\exp_{x}^{M}(t\nu_{x})$ (120)
is an embedding, and there exists a smooth family of metrics $\{g_{t}\}_{t\in[0,\varepsilon]}$ on $\partial M$ such that for any $(x,t)\in\partial M\times[0,\epsilon]$,
$(E^{*}g)_{(x,t)}=(g_{t}\oplus d\tau^{2})_{(x,t)}.$ (121)
Pulling back each metric $g_{t}$ by $\psi$ we have, for any $\xi\in(-4,4)^{m-1}$ and $v,w\in\mathbb{R}^{m-1}$,
$(\psi^{*}g_{t})_{\xi}(v,w)=v^{T}\cdot A_{(\xi,t)}\cdot w,$ (122)
for some positive definite, symmetric $(m-1)$-square matrix $A_{(\xi,t)}$. From the non-degeneracy of the metric and a Lipschitz bound, we have that there exist $C,\tilde{c}>0$ such that for all $t,s\in[0,\epsilon],\xi\in[-3,3]^{m-1},v\in\mathbb{R}^{m}$ we have
$\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)\geq\tilde{c}|v|^{2},$ (123)
$\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)\right|\leq C|v|^{2}|t-s|.$ (124)
Claim 1.
There exists $K>0$ such that for any $t,s\in[0,\epsilon]$, $\xi\in[-3,3]^{m-1},$ $v\in\mathbb{R}^{m}$ such that
$\left|O(\xi,t,s,v)\right|\leq K|v|\cdot|t-s|,$ (125)
where
$O(\xi,t,s,v)=\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v).$ (126)
Proof.
Consider $\tilde{c}>0$ given by (123). We know that there exists $M>0$ such that the map $\sqrt{\cdot}:[\tilde{c},+\infty)\rightarrow\mathbb{R}$ is $M$-Lipschitz. By homogeneity in $|v|$ of (125), we can assume that $|v|=1$. Then the Lipschitz condition and (124) yield
$\left|O(\xi,t,s,v)\right|$ $=\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)\right|$ (127)
$\leq M\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}(v,v)\right|$ (128)
$\leq MC|v|^{2}|t-s|=MC|t-s|.$ (129)
Step 2.
Let $T(\epsilon)\coloneqq E(\partial M\times[0,\epsilon])$ be the $\epsilon$ tubular neighborhood of the boundary. Then $E$ is a diffeomorphism between $\partial M\times[0,\epsilon]$ and $T(\epsilon)$. For fixed $s\in[0,\epsilon]$, consider the product metric $g_{s}\oplus d\tau^{2}$ in $\partial M\times[0,\epsilon]$ and define the metric in $T(\epsilon)$
$\eta_{s}\coloneqq(E^{-1})^{*}(g_{s}\oplus d\tau^{2}).$ (130)
Let $\mathsf{d}_{s}$ be the distance induced from this metric. Consider the map $\Phi:(-4,4)^{m-1}\times[0,\epsilon]\rightarrow M$ given by
$\Phi(\xi,t)=E(\psi(\xi),t)$ (131)
and note that (121) implies that for any $(\xi,t)\in(-4,4)^{m-1}\times[0,\epsilon]$,
$(\Phi^{*}g)_{(\xi,t)}=(\psi^{*}g_{t}\oplus\mathop{}\!\mathrm{d}\tau^{2})_{(\xi,t)}.$ (132)
Set $\mathbb{H}^{m}\coloneqq\{v\in\mathbb{R}^{m}:v_{m}\geq 0\}$. Let $L(\tilde{\gamma})$ be the Euclidean length of a curve $\tilde{\gamma}:[0,1]\to\mathbb{R}^{m}$.
Claim 2.
Let $\tilde{c}>0$ be given by (123). For any $r>0$ and $s\in[0,\varepsilon]$, if a couple $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon]$ is such that $\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-3,3]^{m-1}\times[0,\epsilon]$, then the following holds for any $y\in M$.
  1. (1)
    If $\mathsf{d}(\Phi(\xi,t),y)<r$, then the image of any $\mathsf{d}$-minimizing geodesic $\gamma:[0,1]\rightarrow M$ is contained in $\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$ and $\tilde{\gamma}=\Phi^{-1}\circ\gamma$ satisfies $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$.
  2. (2)
    If $\mathsf{d}_{s}(\Phi(\xi,t),y)<r$, then the image of any $\mathsf{d}_{s}$-minimizing geodesic $\gamma:[0,1]\rightarrow M$ is contained in $\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$ and $\tilde{\gamma}=\Phi^{-1}\circ\gamma$ satisfies $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$.
Proof.
We prove the first result only since the proof of the second one follows from similar lines. Consider a $\mathsf{d}$-minimizing geodesic $\gamma:[0,1]\rightarrow M$ from $\Phi(\xi,t)$ to $y$. Set
$\delta$ $\coloneqq\sup\left\{t\in[0,1]\,:\,\gamma(s)\in\Phi(\mathbb{B}^{m}_{r/\sqrt{c}}(\xi,t)\cap\mathbb{H}^{m})\text{ for any }s\in[0,t)\right\},$
$\tilde{\gamma}$ $\coloneqq\Phi^{-1}\circ\left.\gamma\right|_{[0,\delta]},$
and observe that $\gamma([0,1])\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{c}}(\xi,t)\cap\mathbb{H}^{m})$ if and only if $\delta=1$. We claim that
$L(\tilde{\gamma})\leq\frac{\mathsf{d}_{g}(\Phi(\xi,t),y)}{\sqrt{\tilde{c}}}\,\cdot$ (133)
Indeed, setting $(\tilde{\alpha},\tilde{\gamma}_{m})\coloneqq\tilde{\gamma}$ where $\tilde{\alpha}:[0,\delta]\rightarrow[-3,3]^{m-1}$ and $\tilde{\gamma}_{m}:[0,\delta]\rightarrow[0,\epsilon]$, we have
$\mathsf{d}(\Phi(\xi,t),y)$ $=\int_{0}^{1}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (134)
$\geq\int_{0}^{\delta}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (135)
$=\int_{0}^{\delta}(\Phi^{*}g)_{\tilde{\gamma}(w)}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w$ (136)
[…]
$=\sqrt{\tilde{c}}\,L(\tilde{\gamma}).$ (139)
Now we claim that:
$\delta<1\qquad\Rightarrow\qquad\mathsf{d}_{g}(\Phi(\xi,t),y)\geq r.$ (140)
Indeed, if $\delta<1$, since $\tilde{\gamma}(0)=(\xi,t)$ and $\tilde{\gamma}(\delta)\in\partial\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}$, then
$L(\tilde{\gamma})\geq\frac{r}{\sqrt{\tilde{c}}}$ (141)
and we get $\mathsf{d}_{g}(\Phi(\xi,t),y)\geq r$ from (133). Therefore, if $\mathsf{d}_{g}(\Phi(\xi,t),y)<r$, then (140) implies that $\delta=1$ which means $\gamma([0,1])$ is included in $\Phi(\mathbb{B}^{m}_{r/\sqrt{c}}(\xi,t)\cap\mathbb{H}^{m})$, and (133) yields $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$ as desired. ∎
Claim 3.
There exists $K>0$ such that for all $(\xi,t),(\eta,s)\in[-2,2]^{m-1}\times[0,\epsilon]$ and $r>0$ such that $\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-3,3]^{m-1}\times[0,\epsilon]$:
$\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))<r\quad\Rightarrow\quad\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))<r+Kr^{2},$ (142)
$\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))\geq r\quad\Rightarrow\quad\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))\geq r-Kr^{2}.$ (143)
Proof.
Suppose that $\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))<r$. Let $\gamma:[0,1]\rightarrow M$ be a $\mathsf{d}$-minimizing geodesic between $\Phi(\xi,t)$ and $\Phi(\eta,s)$. Then by Claim 2, we have that $\gamma([0,1])\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$, and by defining $(\tilde{\alpha},\tilde{\gamma}_{m})=\tilde{\gamma}\coloneqq\Phi^{-1}\circ\gamma$, we have that $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$. Thus we obtain
$\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))$ $=\int_{0}^{1}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (144)
$=\int_{0}^{1}g_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w-\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (145)
$\quad+\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (146)
$=\int_{0}^{1}\left(\psi^{*}g_{\tilde{\gamma}_{m}(w)}\oplus d\tau^{2}\right)_{(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w))}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w$ (147)
$\quad-\int_{0}^{1}\left(\psi^{*}g_{t}\oplus d\tau^{2}\right)_{(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w))}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w$ (148)
$\quad+\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (149)
$=\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (150)
$\quad+\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w,$ (151)
where we use (126) to get the last equality. By Claim 1, we have that
$\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w\right|$ $\leq\int_{0}^{1}\left|O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\right|\mathop{}\!\mathrm{d}w$ (152)
$\leq\int_{0}^{1}K\left|\dot{\tilde{\gamma}}(w)\right|\left|t-\tilde{\gamma}_{m}(w)\right|\mathop{}\!\mathrm{d}w.$ (153)
By Claim 2, we have that $|t-\tilde{\gamma}_{m}(s)|<\frac{r}{\sqrt{\tilde{c}}}$ and $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$, hence we get
$\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w\right|$ $\leq K\frac{r}{\sqrt{\tilde{c}}}\int_{0}^{1}|\dot{\tilde{\gamma}}(w)|\mathop{}\!\mathrm{d}w<K\frac{r^{2}}{\tilde{c}}\,\cdot$ (154)
Thus
$\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))$ $\leq\int_{0}^{1}\left(\eta_{t}\right)_{\gamma(w)}^{\frac{1}{2}}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (155)
$\leq\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))+\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\gamma}(w))\mathop{}\!\mathrm{d}w\right|$ (156)
$\leq\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))+K\frac{r^{2}}{\tilde{c}}$ (157)
$<r+K\frac{r^{2}}{\tilde{c}}\,\cdot$ (158)
where we use (151) to get the second inequality and (154) to get the third one. This proves (142).
To prove (143), we may assume $\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))\geq r$ and $\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))<r$. Let $\gamma:[0,1]\rightarrow M$ be a geodesic in the metric $g_{t}\oplus d\tau^{2}$. By Claim 2, we have $\gamma([0,1])\subset\Phi(\mathbb{B}^{m}_{\frac{r}{\sqrt{\tilde{c}}}}(\xi,t)\cap\mathbb{H}^{m})$ with $(\tilde{\alpha},\tilde{\gamma}_{m})=\tilde{\gamma}=\Phi^{-1}\circ\gamma$ satisfying $L(\tilde{\gamma})<\frac{r}{\sqrt{\tilde{c}}}$. The same estimates as before are satisfied, hence we conclude:
$r$ $<\mathsf{d}(\Phi(\xi,t),\Phi(\eta,s))\leq\int_{0}^{1}g^{\frac{1}{2}}_{\gamma(w)}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w$ (159)
$\leq\int_{0}^{1}(\eta_{t})^{\frac{1}{2}}_{\gamma(w)}(\dot{\gamma}(w),\dot{\gamma}(w))\mathop{}\!\mathrm{d}w+\left|\int_{0}^{1}O(\tilde{\alpha}(w),\tilde{\gamma}_{m}(w),t,\dot{\tilde{\gamma}}(w))\mathop{}\!\mathrm{d}w\right|$ (160)
$\leq\mathsf{d}_{t}(\Phi(\xi,t),\Phi(\eta,s))+\frac{K}{\tilde{c}}r^{2}.$ (161)
We omit the proof of the next elementary claim.
Claim 4.
Let $(x,t),(y,\tau)\in\partial M\times[0,\epsilon]$ and $s\in[0,\epsilon]$. Let $\gamma:[0,1]\rightarrow\partial M$ be the geodesic between $x$ and $y$ in the metric $g_{s}$. Then the curve
$w\in[0,1]\mapsto(\gamma(w),(1-w)t+w\tau)$ (162)
is a geodesic in the metric $g_{s}\oplus d\tau^{2}$.
Step 3.
For every $t\in[0,\epsilon]$, consider the exponential map $\exp^{g_{t}}$ given by the metric $g_{t}$. Since $g_{t}$ varies smoothly with respect to $t$, and $\partial M\times[0,\epsilon]$ is compact, we know that there exists $\delta>0$ lower than the injectivity radius of each $\exp^{g_{t}}$. For any $(\xi,t,\zeta,s)$ in
$D_{\delta}:=[-1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$
define
$\Psi(\xi,t,\zeta,s)\coloneqq E(\exp^{g_{t}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta),s).$ (163)
We may write
$\Psi_{(\xi,t)}(\zeta,s)=\Psi(\xi,t,\zeta,s)$
to see $\Psi$ as a function of the two last variables only, the two first being frozen. Observe that for any $(\xi,t,\zeta)$ in $[-1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)$,
$(g_{t})^{\frac{1}{2}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta,(d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ $=(\psi^{*}g_{t})_{\xi}^{\frac{1}{2}}(A_{(\xi,t)}^{-\frac{1}{2}}\zeta,A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ (164)
[…]
$=|\zeta|<\delta.$ (166)
Since $\delta$ is lower than the injectivity radius of the exponentials, the map
$\zeta\in\mathbb{B}^{m-1}_{\delta}(0)\mapsto\exp^{g_{t}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ (167)
is injective. Thus for every $(\xi,t)\in[-1,1]^{m-1}\times[0,\epsilon]$ the map $\Psi_{(\xi,t)}$ defined on $\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$ is a local parametrization of $M$. Moreover,
$\Psi_{(\xi,t)}(0,t)=E(\psi(\xi),t)=\Phi(\xi,t),$ (168)
and
$\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(0,t)}\right)=1.$ (169)
Claim 5.
Consider $(\xi,t)\in[-1,1]^{m-1}\times[0,\epsilon]$, and $(\zeta,s)\in\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$. Then
$\mathsf{d}_{t}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))=\sqrt{|\zeta|^{2}+(t-s)^{2}}.$ (170)
Proof.
By Claim 4, we know that the geodesic between $E^{-1}(\Psi_{(\xi,t)}(0,t))$ and $E^{-1}(\Psi_{(\xi,t)}(\zeta,s))$ in the metric $g_{s}\oplus d\tau^{2}$ is
$\gamma:w\in[0,1]$ $\mapsto(\exp^{g_{t}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}w\zeta),(1-w)t+ws)$
$\eqqcolon(\tilde{\gamma}(w),\gamma_{m}(w)).$
Then we also know that, for any $w\in[0,1]$,
$(g_{t})_{\tilde{\gamma}(w)}^{\frac{1}{2}}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))=|\zeta|.$ (171)
Thus
$\mathsf{d}_{t}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))$ $=\int_{0}^{1}\sqrt{(g_{t})_{\tilde{\gamma}(w)}^{2}(\dot{\tilde{\gamma}}(w),\dot{\tilde{\gamma}}(w))+(s-t)^{2}}\mathop{}\!\mathrm{d}w$ (172)
$=\sqrt{|\zeta|^{2}+(s-t)^{2}}.$ (173)
Claim 6.
There exist $r_{0},\kappa>0$ such that for all $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$ and $r\in(0,r_{0})$ such that $\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\subset[-3,3]^{m-1}\times[0,\epsilon]$, we have
$\mathcal{L}^{m}\left(\mathbb{B}^{m}_{r}(0,t)\,\triangle\,\bigg{(}\Psi_{(\xi,t)}^{-1}\left(B_{r}(\Psi_{(\xi,t)}(0,t))\right)\right)\bigg{)}\leq\kappa r^{m+1}.$ (174)
Proof.
For any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, there exists $r_{0}(\xi,t)>0$ small enough such that
$B_{r_{0}}(\Psi_{(\xi,t)}(0,t))=B_{r_{0}}(\Phi(\xi,t))\subset\Psi_{(\xi,t)}(\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]).$ (175)
By compactness of $[-2,2]^{m-1}\times[0,\epsilon/2]$ and continuity of the maps $\Psi_{(\xi,t)}^{-1}$, we get that there exists a common $r_{0}>0$ such that the previous holds for any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$. Consider $r<r_{0}$ and $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, then
$B_{r}(\Psi_{(\xi,t)}(0,t))\subset\text{Im}(\Psi_{(\xi,t)}).$ (176)
Set $A_{1}:=\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))$ and $A_{2}:=\mathbb{B}^{m}_{r}(0,t).$ We will show that there exists $K>0$ such that
$A_{1}\backslash A_{2}\subset\mathbb{B}^{m}_{r+Kr^{2}}(0,t)\backslash\mathbb{B}^{m}_{r}(0,t).$ (177)
For $(\zeta,s)\in A_{1}\backslash A_{2}$, we know that
$\mathsf{d}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))<r.$ (178)
Therefore, from (168) and Claim 3, we conclude that there exists $K>0$ such that
$\mathsf{d}_{t}(\Psi_{(\xi,t)}(\zeta,s),\Psi_{(\xi,t)}(0,t))<r+Kr^{2}.$ (179)
Then we get from Claim 5 that
$\sqrt{|\zeta|^{2}+(t-s)^{2}}<r+Kr^{2}$ (180)
hence (177) is proved. A similar proof shows that
$A_{2}\backslash A_{1}\subset\mathbb{B}^{m}_{r}(0,t)\backslash\mathbb{B}^{m}_{r-Kr^{2}}(0,t).$
From the latter and (177) we conclude that
$A_{1}\triangle A_{2}\subset\mathbb{B}^{m}_{r+Kr^{2}}(0,t)\backslash\mathbb{B}^{m}_{r-Kr^{2}}(0,t).$ (181)
Thus
$\mathcal{L}^{m}(A_{1}\triangle A_{2})$ $\leq\mathcal{L}^{m}(\mathbb{B}^{m}_{r+Kr^{2}}(0,t)\backslash\mathbb{B}^{m}_{r-Kr^{2}}(0,t))$ (182)
$\leq r^{m}\mathcal{L}^{m}(\mathbb{B}^{m}_{1+Kr}(0)\backslash\mathbb{B}^{m}_{1-Kr}(0))$ (183)
$=r^{m}\omega_{m}((1+Kr)^{m}-(1-Kr)^{m})$ (184)
$\leq r^{m}C(Kr)\qquad\qquad\qquad\qquad\qquad\qquad\text{for some$C>0$}$ (185)
$\leq(CK)r^{m+1}.$ (186)
Claim 7.
There exist $C,r_{0}>0$ such that for all $r\in(0,r_{0})$ and
$(\xi,t,\zeta,s)\in[-1,1]^{m-1}\times[0,\epsilon/4]\times\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]$
such that $\mathbb{B}^{m}_{2r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-2,2]^{m-1}\times[0,\epsilon/2]$ and $\mathsf{d}(\Psi_{(\xi,t)}(\zeta,s),\Phi(\xi,t))<r$, then
$\left|\frac{1}{V(\Psi_{(\xi,t)}(\zeta,s),r)}-\frac{1}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\right|\leq\frac{C}{r^{m-1}}\,\cdot$ (187)
Proof.
Let us first consider the map $G:D_{\delta}\rightarrow\mathbb{R}$ given by
$G(\xi,t,\zeta,s)\coloneqq\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(\zeta,s)}\right)^{\frac{1}{2}}.$ (188)
This map is $C^{\infty}$, and its value at any $(\xi,t)\times(0,t)$ is $1$, thus by a Taylor expansion in the variable $(\zeta,s)$ centered at $(0,t)$, and compactness of $D_{\delta}$, we obtain that there exists $k>0$ such that for all $(\xi,t,\zeta,s)\in D_{\delta}$ we have
$\left|\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(\zeta,s)}\right)^{\frac{1}{2}}-1\right|\leq k|(\zeta,s)-(0,t)|.$ (189)
In particular there exists $C>0$ such that for all $(\xi,t,\zeta,s)\in D_{\delta}$,
$\left|\det\left(\left[\Psi_{(\xi,t)}^{*}g\right]_{(\zeta,s)}\right)^{\frac{1}{2}}\right|\leq C.$ (190)
Let us now consider $r,\xi,t,\zeta,s$ as in the statement of the claim. Set $y\coloneqq\Psi_{(\xi,t)}(\zeta,s)$. Since $\mathsf{d}(y,\Phi(\xi,t))<r$, and $\mathbb{B}^{m}_{2r/\sqrt{\tilde{c}}}(\xi,t)\subset[-2,2]^{m-1}\times[0,\epsilon/2]$, we know by Claim 2 that $y\in\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m})$. Thus if we set
$(\eta,\tau)\coloneqq\Phi^{-1}(y)\in[-2,2]^{m-1}\times[0,\epsilon/2],$ (191)
we obtain that
$\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\eta,\tau)\cap\mathbb{H}^{m}\subset\mathbb{B}^{m}_{2r/\sqrt{\tilde{c}}}(\xi,t)\cap\mathbb{H}^{m}\subset[-2,2]^{m-1}\times[0,\epsilon/2].$ (192)
Thus by Claim 2, we conclude that $B_{r}(y)=B_{r}(\Phi(\eta,\tau))\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\eta,\tau))$.
By (168) and (131), we easily see that
$s=\tau.$ (193)
Moreover, by (191) and (193), we also have that
$\Psi_{(\eta,s)}(0,s)=E(\psi(\eta),s)=\Phi(\eta,s)=y.$ (194)
Choose $r_{0}$ such that $\frac{r_{0}}{\sqrt{\tilde{c}}}<\frac{\epsilon}{4}$. Since $t\leq\frac{\epsilon}{4}$ and $(\eta,s)\in\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t)$, this implies that $s\leq\frac{\epsilon}{2}$. Thus we can use Claim 6 to ensure that
$\mathcal{L}^{m}\left(\mathbb{B}^{m}_{r}(0,s)\triangle\left(\Psi_{(\eta,s)}\right)^{-1}\left(B_{r}(\Psi_{(\eta,s)}(0,s))\right)\right)\leq\kappa r^{m+1}.$ (195)
Then
$V(y,r)$ $=\int_{\Psi_{(\eta,z)}^{-1}(B_{r}(y))}\left|\det\left[\Psi_{(\eta,s)}^{*}g\right]_{w}\right|^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(w)$ (196)
[…]
that is, there exists $\tilde{C}>0$ such that
$\left|\mathrm{vol}_{g}(B_{r}({y}))-\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})\right|\leq\tilde{C}r^{m+1}$ (199)
Thus using the local Ahlfors regularity of $(M,g)$ and Claim 6, we obtain
$\left|\frac{1}{\mathrm{vol}_{g}(B_{r}(\overline{y}))}-\frac{1}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})}\right|$ $=\left|\frac{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})-\mathrm{vol}_{g}(B_{r}(\overline{y}))}{\mathrm{vol}_{g}(B_{r}(\overline{y}))\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})}\right|$ (200)
$\leq\frac{C\tilde{C}r^{m+1}}{r^{m}\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,z)\cap\mathbb{H}^{m})}$ (201)
$\leq\frac{C\tilde{C}}{\mathcal{L}^{m}(\mathbb{B}^{m}_{1}(0)\cap\mathbb{H}^{m})r^{m-1}}$ (202)
concluding the proof. ∎
Step 4.
Claim 8.
For all $f\in C^{\infty}(M)$, there exists $C>0$ such that for all $(\xi,t,\zeta,s)\in D_{\delta}$ we have
$\left|f\circ\Psi_{(\xi,t)}(\zeta,s)-f\circ\Psi_{(\xi,t)}(0,t)\right|\leq C\|(\zeta,s)-(0,t)\|_{2},$ (203)
$\left|f\circ\Psi_{(\xi,t)}(\zeta,s)-f\circ\Psi_{(\xi,t)}(0,t)-\nabla(f\circ\Psi_{(\xi,t)})_{(0,t)}\cdot\left((\zeta,s)-(0,t)\right)\right|\leq C\|(\zeta,s)-(0,t)\|_{2}^{2}$ (204)
Also if $\partial_{\nu}f|_{\partial M}=0$, then there exists $C>0$ such that
$\left|\partial_{m}f\circ\Psi_{(\xi,t)}(0,t)\right|\leq Ct$ (205)
Proof.
The map $\tilde{f}((\xi,t),(\zeta,s))=f\circ\Psi_{(\xi,t)}(\zeta,s)$ is $C^{\infty}$, thus by a Taylor expansion of order $1$ and $2$ respectively, and compactness of $D_{\delta}$, we conclude the first two inequalities. For the last, we notice that
$\partial_{m}f\circ\Psi_{(\xi,0)}(0,0)=(\partial_{\nu}f)(\psi(\xi,0))=0.$ (206)
Thus by a Taylor expansion and compactness of $D_{\delta}$ we conclude (205). ∎
Now we will fix a function $f\in C^{\infty}(M)$ such that $\partial_{\nu}f|_{\partial M}=0$, and show that for $x\in M$ such that $d(x,\partial M)<\frac{\epsilon}{4}$ then $\Delta_{r}f(x)$ is uniformly bounded. The proof for points $x$ with $d(x,\partial M)\geq\frac{\epsilon}{4}$, follows from the uniform convergence obtained in Proposition 5.3. We will study the following term of the AMV:
$G_{r}(x):=\frac{1}{r^{2}}\int_{B_{r}(x)}\frac{1}{\mathrm{vol}_{g}(B_{r}(y))}\left(f(y)-f(x)\right)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y)$ (207)
since the bound for the remainder follows similarly.
We notice that by equation (119) we have that there exists some $i\in\{1,...,l\}$ and $(\xi,t)\in[-1,1]^{m-1}\times[0,\frac{\epsilon}{4}]$ such that $x=\Phi_{i}(\xi,t)$. We let $\Phi=\Phi_{i}$. Also by (175), for $0<r<r_{0}$ in the conditions of the Claim, we have that $B_{r}(x)=B_{r}(\Psi_{(\xi,t)}(0,t))\subset\Psi_{(\xi,t)}(\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon])$. Also we can choose $r_{0}$ small enough so that for all $(\xi,t)\in[-1,1]^{m-1}\times[0,{\epsilon}/{4}]$ we have $\mathbb{B}_{2r_{0}/\sqrt{\tilde{c}}}(x,t)\cap\mathbb{H}^{m}\subset[-2,2]^{m-1}\times[0,\epsilon/2]$. With this we can apply Claim 2 to conclude that $B_{r}(\Psi_{(\xi,t)(0,t)})\subset\Phi(\mathbb{B}^{m}_{r/\sqrt{\tilde{c}}}(\xi,t))$ and we can also apply Claim 7 for points $\Psi_{(\xi,t)}(\zeta,s)\in B_{r}(\Psi_{(\xi,t)}(0,t))$.
Thus we can change variables of the integral to obtain
$G_{r}(x)=\frac{1}{r^{2}}\int_{\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))}\frac{1}{\mathrm{vol}_{g}(B_{r}(\Psi_{(\xi,t)}(\zeta,s)))}$ (208)
×$\times\left(f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))\right)\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)$ (209)
$\begin{split}G_{r}(x)&=\frac{1}{r^{2}}\int_{\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 7 and 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 6 and 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by LABEL:lem:determinantLemma and 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\sum_{i=1}^{m-1}\partial_{j}(f\circ\Psi_{(\xi,t)})(0,t)\zeta_{i}+\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\sum_{i=1}^{m-1}\partial_{j}(f\circ\Psi_{(\xi,t)})(0,t)\zeta_{i}+\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s))+O(1)\end{split}$ (by 8)
$\begin{split}&=\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s))+O(1)\end{split}$ (by symmetry)
Now we separate further in two cases. First if $\mathsf{d}(x,\partial M)>2r$ then $t>2r$ and so for all $(\zeta,s)\in\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}$ we have $\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})=\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0))$, and so we conclude that
$\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s)$ (211)
$=\frac{1}{r^{2}\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0))}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\partial_{m}(f\circ\Psi_{(\xi,t)})(0,t)(s-t)\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(\zeta)\mathop{}\!\mathrm{d}\mathcal{L}(s)$ (212)
$=0.$ (213)
This shows that $G_{r}(x)=O(1)$ for $x$ such that $2r<d(x,\partial M)<\frac{\epsilon}{4}$ since there are only a finite number of parametrizations. On the other hand if $d(x,\partial M)\leq 2r$, then we have by Lemma 8 that for $t\leq 2r$ then $\left|\partial_{m}(f\circ\Psi_{(x,t)})(0,t)\right|\leq 2Cr$, and so
$\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(x,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(y)\mathop{}\!\mathrm{d}\mathcal{L}(s)=O(1),$ (214)
which shows that $G_{r}(x)=O(1)$ for $x$ such that $d(x,\partial M)<2r$.
6. Spectral convergence
In this section, we prove Theorem 3. We consider a smooth, compact, connected manifold $M^{m}$ endowed with a smooth Riemannian metric $g$. We let $\mathsf{d}_{g}$ and $\mathrm{vol}_{g}$ be the associated Riemannian distance and volume measure on $M$, respectively. If $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$), we let $\{\mu_{k}\}_{k\in\mathbb{N}}$ be the sequence of Laplace (resp. Neumann) eigenvalues of $(M,g)$.
6.1. Existence of limit eigenfunctions
Recalling that $C^{\infty}_{\nu}(M)$ is defined in (5), we define the Hilbert space
$H\coloneqq\overline{C^{\infty}_{\nu}(M)}^{\|\cdot\|_{W^{2,2}}}.$
We let $\Pi(M,\mathrm{vol}_{g})$ be defined as in (42) and we consider the operator $T:\Pi(M,\mathrm{vol}_{g})\rightarrow H^{*}$ which maps any $f\in\Pi(M,\mathrm{vol}_{g})$ to
$T(f)\coloneqq\left(v\in H\mapsto-\int_{M}f\Delta_{g}v\mathop{}\!\mathrm{d}\mathrm{vol}_{g}\right).$ (215)
Lemma 6.1.
The operator $T$ is injective.
Proof.
For $f\in\Pi(M,\mathrm{vol}_{g})$ such that $f\neq 0$, let $v\in H$ be the solution of
$\begin{cases}-{\Delta}_{g}v=f&\text{in }M,\\ \partial_{\nu}v=0&\text{in }\partial M.\end{cases}$ (216)
The solution to this problem exists since $\int_{M}f=0=\int_{\partial M}\partial_{\nu}v$. In fact by regularity theory we can conclude that $v\in H$, and so we conclude that
$T(f)(v)=-\int_{M}f{\Delta}_{g}v\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\int_{M}f^{2}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}\neq 0.$ (217)
Thus $T(f)\neq 0$, concluding that $T$ is injective. ∎
Let us now prove the existence of $L^{2}$-weak limit eigenfunctions.
Proposition 6.2.
Let $(r_{n})$ be a sequence of positive numbers such that $r_{n}\to 0$. For any $n$ let $(\lambda_{k,{r_{n}}})$ be the eigenvalues of the operator $\tilde{\Delta}_{r_{n}}$ and let $(f_{k,r_{n}})$ be corresponding eigenfunctions. Then for any $k$ there exists a Laplace (resp. Neumann) eigenfunction $f$ of $(M,g)$ with associated eigenvalue $\mu$ such that, up to extracting subsequences, satisfy
$f_{k,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\rightharpoonup}}f,$ (218)
$\lambda_{k,r_{n}}\to C_{m}\,\mu,$ (219)
$\sup_{n}E_{r_{n}}(f_{k,r_{n}}-f)<+\infty.$ (220)
Proof.
By the proof of Theorem 1, in particular (54), there exists $\lambda\geq 0$ and a subsequence such that $\lambda_{k}(-\tilde{\Delta}_{r_{n}})\rightarrow\lambda$ up to subsequence. Since $\|f_{k,r_{n}}\|_{L^{2}(M)}=1$ for any $n$, there exists $f\in L^{2}(M,\mathrm{vol}_{g})$ such that the weak convergence (218) holds up to subsequence. Therefore, by Theorem 2, we get that for any $\psi\in C^{\infty}(M)$ (resp. $C^{\infty}_{\partial_{\nu}}(M)$),
$\int_{M}f\Delta_{g}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}$ $=\lim_{n}\frac{1}{C_{m}}\int_{M}f_{k,r_{n}}\tilde{\Delta}_{r_{n}}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\frac{1}{C_{m}}\lim_{n}\int_{M}\left(\tilde{\Delta}_{r_{n}}f_{k,r_{n}}\right)\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}$
$=-\frac{1}{C_{m}}\lim_{n}\lambda_{k}(-\tilde{\Delta}_{r_{n}})\int_{M}f_{k,r_{n}}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=-\frac{1}{C_{m}}\lambda\int_{M}f\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}.$
Moreover, since for any $n$ it holds that $\lambda_{k,r_{n}}>0$ and
$0=\int_{M}-\tilde{\Delta}_{r_{n}}f_{k,r_{n}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\int_{M}\tilde{\lambda}_{k,r_{n}}f_{k,r_{n}}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}$
we get that $\int_{M}f_{k,r_{n}}=0$. Thus $\int_{M}f\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=0$ by weak convergence.
Now let $v\in W^{2,2}(M)$ be the solution of
$\begin{cases}-{\Delta}_{g}v=\frac{\lambda}{C_{m}}f&\text{ in }M\\ \partial_{\nu}v=0&\text{ in }\partial M\end{cases}$ (221)
satisfying $\int_{M}v=0$. Then we have that for $\psi\in H\cap C^{\infty}(M)$ then
$\int_{M}v{\Delta}_{g}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=-\frac{1}{C_{m}}\lambda\int_{M}f\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}=\int_{M}f{\Delta}_{g}\psi\mathop{}\!\mathrm{d}\mathrm{vol}_{g}.$ (222)
Since this is a dense subspace of $H$ and the functionals are continuous with respect to $W^{2,2}(M)$ in $\psi$, the equality holds for all $H$. Thus by Lemma 6.1 we conclude that $v=f$, and so $f$ satisfies
$\begin{cases}-{\Delta}_{g}f=\frac{\lambda}{C_{m}}f&\text{ in }M\\ \partial_{\nu}f=0&\text{ in }\partial M\end{cases}$ (223)
thus $f$ is a Neumann eigenfunction, and so it must be $C^{\infty}(M)$. Also since both $f_{k,r_{n}},f\in\Pi(M,\mathrm{vol}_{g})$, we know by Proposition 3.9 using triangle inequality of the inner product,
$E_{r_{n}}(f_{k,r_{n}}-f)^{\frac{1}{2}}\leq E_{r_{n}}(f_{k,r_{n}})^{\frac{1}{2}}+E_{r_{n}}(f)^{\frac{1}{2}}.$ (224)
We know that $E_{r_{n}}(f_{k,r_{n}})=\lambda_{k,r_{n}}\|f_{k,r_{n}}\|_{L^{2}(M)}=\lambda_{k,r_{n}}$ which is uniformly bounded. Also since $f\in C^{\infty}(M)$, by Lemma 5.1, we know that $E_{r_{n}}(f)$ is also uniformly bounded, concluding the proof. ∎
6.2. Energy comparison
Let us now compare the energy of a map defined on $M$ with the energy of the image of the map through a local chart parametrizing a neighborhood of an open subset of $\partial M$. To this aim, up to scaling, we consider a map $\Phi:(-1,1)^{m-1}\times[0,1)\to M$ which is a bi-Lipschitz homeomorphism onto its image. We set
$\mathcal{Q}\coloneqq(-1/2,1/2)^{m-1}\times[0,1/2).$ (225)
Lemma 6.3.
There exist constants $\tilde{C}=\tilde{C}(\Phi)>0$ and $\tilde{c}=\tilde{c}(\Phi)>0$ such that for any $f\in L^{2}(M)$, for any $r\in(0,1/2)$,
$\tilde{E}_{\tilde{c}r,\mathfrak{Q}}(f\circ\Phi)\leq\tilde{C}E_{r,\mathfrak{M}}(f).$ (226)
where $\mathfrak{Q}\coloneqq(\mathcal{Q},d_{\infty},\mathcal{L}^{m})$ and $\mathfrak{M}\coloneqq(M,d,\mu)$.
Proof.
We start by pointing out that there exist constants $c=c(\Phi)>0$ and $C=C(\Phi)>0$ such that for all $x\in\mathcal{Q}$ and $r\in(0,1/2)$,
$\Phi(Q_{cr}(x))\subset B_{r}(\Phi(x))\subset\Phi(Q_{Cr}(x)),$
$V(\Phi(x),r)\leq C\mathcal{L}^{m}(Q_{cr}(x)\cap\mathcal{Q}),$
$\det(g_{x})\geq 0,$
where $g_{x}$ is the metric in the coordinates given by $\Phi$. Then for any $x,y\in\mathcal{Q}$,
$\tilde{a}_{r,\mathfrak{M}}(\Phi(x),\Phi(y))$ $=1_{B_{r}(\Phi(x))}(\Phi(y))\left(\frac{1}{V(\Phi(x),r)}+\frac{1}{V(\Phi(y),r)}\right)$
$\geq 1_{Q_{cr}(x)}(y)\left(\frac{1}{C\mathcal{L}^{m}(Q_{cr}(x)\cap\mathcal{Q})}+\frac{1}{C\mathcal{L}^{m}(Q_{cr}(y)\cap\mathcal{Q})}\right)$
$=\frac{\tilde{a}_{cr,\mathfrak{Q}}(x,y)}{C}\,\cdot$
Thus
$\tilde{E}_{r,\mathfrak{M}}(f)$ $\geq\iint_{\Phi(\mathcal{Q})^{2}}\tilde{a}_{r,\mathfrak{M}}(p,q)\left(\frac{f(p)-f(q)}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(q)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(p)$
$=\iint_{\mathcal{Q}^{2}}\tilde{a}_{r,\mathfrak{M}}(\Phi(x),\Phi(y))\left(\frac{f(\Phi(x))-f(\Phi(y))}{r}\right)^{2}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\sqrt{\det(g_{x})\det(g_{y})}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
$\geq\iint_{\mathcal{Q}^{2}}\frac{c}{C}\tilde{a}_{cr,\mathfrak{Q}}(x,y)\left(\frac{f(\Phi(x))-f(\Phi(y))}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
$=\frac{c}{C}E_{cr,\mathfrak{Q}}(f\circ\Phi).$
Taking $\tilde{c}=c$ and $\tilde{C}=\frac{C}{c}$, we obtain the result. ∎
6.3. Proof of Theorem 3
We are now in a position to prove Theorem 3. Recall the context of this result : $(r_{n})\subset(0,+\infty)$ is a sequence such that $r_{n}\to 0$, $(M^{m},g)$ is a compact, connected, smooth Riemannian manifold with $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$ ), $k$ is a positive integer, $\mu_{k}$ is the $k$-th lowest Laplace (resp. Neumann) eigenvalue of $\Delta_{g}$, and $f_{k,r_{n}}$ is an eigenfunction of $-\tilde{\Delta}_{r_{n}}$ associated with the $k$-th eigenvalue $\lambda_{k}(-\tilde{\Delta}_{r_{n}})$ of this operator.
Proof.
Step 1. First we show strong $L^{2}$-convergence of the sequence $(f_{k,r_{n}})$. We proceed by contradiction. By Proposition 6.2, we can assume that there exist $\alpha>0$, $f\in L^{2}(M,\mu)$ which is a Neumann eigenfunction, and $(r_{n})\subset(0,+\infty)$ such that $r_{n}\to 0$ and
$f_{k,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\rightharpoonup}}f,\quad\quad\|f_{k,r_{n}}-f\|_{L^{2}(M)}^{2}\geq\alpha.$ (227)
Since $M$ is a compact manifold with boundary, up to scaling there exist finitely many bi-Lipschitz homeomorphisms $\{\Phi_{j}:(-1,1)^{m-1}\times[0,1)\rightarrow M\}_{j\in\{1,\ldots\ell\}}$ such that
$\bigcup_{j}\Phi_{j}(\mathcal{Q})=M,$ (228)
where $\mathcal{Q}$ is as in (225), and
$\bigcup_{j}\Phi_{j}\Big{(}(-1/2,1/2)^{m-1}\times\{0\}\Big{)}=\partial M.$ (229)
Then there exists $j\in\{1,\ldots,\ell\}$ such that, up to a subsequence,
$\inf_{n}\int_{\Phi_{j}(\mathcal{Q})}|f_{k,r_{n}}-f|^{2}\mathop{}\!\mathrm{d}\mathrm{vol}_{g}\geq\frac{\alpha}{\ell}>0.$ (230)
From this we conclude that there exists $\tilde{\alpha}>0$ such that
$\inf_{n}\int_{\mathcal{Q}}|f_{k,r_{n}}-f|^{2}\circ\Phi_{j}\mathop{}\!\mathrm{d}\mathcal{L}^{m}\geq\tilde{\alpha}>0.$ (231)
Let us set $\Phi:=\Phi_{j}$. Then there exist $C,\tilde{C},\tilde{c}>0$ such that for any $n$,
[…]
By the weak convergence we also have
$h_{n}\coloneqq(f_{k,r_{n}}-f)\circ\Phi\stackrel{{\scriptstyle L^{2}}}{{\rightharpoonup}}0.$ (232)
Let us set $\overline{r}_{n}=\tilde{c}r_{n}$. For an integer $N$ to be chosen later, consider a decomposition of $\mathcal{Q}$ into $L_{N}$ disjoint subcubes $\{\tilde{Q}_{i}\}$ of size $1/N$. For any $x,y\in\mathcal{Q}$, we set
$a_{r}(x,y)\coloneqq\chi_{Q_{r}(x)\cap\mathcal{Q}}(y)\left(\frac{1}{\mathcal{L}^{m}(Q_{r}(x)\cap\mathcal{Q})}+\frac{1}{\mathcal{L}^{m}(Q_{r}(y)\cap\mathcal{Q})}\right),$ (233)
$a_{r,i}(x,y)\coloneqq\chi_{Q_{r}(x)\cap\mathcal{Q}_{i}}(y)\left(\frac{1}{\mathcal{L}^{m}(Q_{r}(x)\cap\mathcal{Q}_{i})}+\frac{1}{\mathcal{L}^{m}(Q_{r}(y)\cap\mathcal{Q}_{i})}\right),$ (234)
and we point out that for $x,y\in\mathcal{Q}_{i}$
$a_{r}(x,y)\geq\frac{1}{2^{m}}a_{r,i}(x,y).$ (235)
We also set for any $n$,
$\epsilon_{i,n,N}\coloneqq\int_{\mathcal{Q}_{i}}h_{n}\mathop{}\!\mathrm{d}\mathcal{L}^{m},\qquad\delta_{n,N}\coloneqq\max_{i}|\epsilon_{i,n,N}|.$ (236)
We obtain that for any $n$,
$\tilde{E}_{\overline{r}_{n},\mathfrak{Q}}(h_{n})$ $=\int_{\mathcal{Q}}\left(\int_{\mathcal{Q}}a_{\overline{r}_{n}}(x,y)\frac{\left(h_{n}(x)-h_{n}(y)\right)^{2}}{\overline{r}_{n}^{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (237)
$=\sum_{i}\int_{\mathcal{Q}_{i}}\left(\int_{\mathcal{Q}}a_{\overline{r}_{n}}(x,y)\frac{\left(h_{n}(x)-h_{n}(y)\right)^{2}}{\overline{r}_{n}^{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (238)
$\geq\frac{1}{2^{m}}\sum_{i}\int_{\mathcal{Q}_{i}}\left(\int_{\mathcal{Q}_{i}}a_{\overline{r}_{n},i}(x,y)\frac{\left(h_{n}(x)-h_{n}(y)\right)^{2}}{\overline{r}_{n}^{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\right)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$ (239)
$=\frac{1}{2^{m}}\sum_{i}\tilde{E}_{\overline{r}_{n},\mathcal{Q}_{i}}(h_{n})$ (240)
$=\frac{1}{2^{m}}\sum_{i}\tilde{E}_{\overline{r}_{n},\mathcal{Q}_{i}}(h_{n}-\epsilon_{i,n,N})$ (241)
$\geq\frac{1}{2^{m}}\sum_{i}\|h_{n}-\epsilon_{i,n,N}\|_{L^{2}(\mathcal{Q}_{i})}^{2}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathcal{Q}_{i}})$ (242)
$\geq\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\sum_{i}\|h_{n}-\epsilon_{i,n,N}\|_{L^{2}(\mathcal{Q}_{i})}^{2}$ (243)
$=\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\sum_{i}\left(\|h_{n}\|_{L^{2}(\mathcal{Q}_{i})}^{2}-2\epsilon_{i,n,N}\int_{\mathcal{Q}_{i}}h_{n}\mathop{}\!\mathrm{d}\mathcal{L}^{m}+\mathcal{L}^{m}(\mathcal{Q}_{i})\epsilon_{i,n,N}^{2}\right)$ (244)
$\geq\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\left(\|h_{n}\|^{2}_{L^{2}(\mathcal{Q})}-3L_{n}\delta_{n,N}\right)$ (245)
$\geq\frac{\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1/N)})}{2^{m}}\left(\tilde{\alpha}-3L_{n}\delta_{n,N}\right).$ (246)
By Lemma 4.3, we choose $N$ big enough to ensure that for any $n$,
$\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}(1/N)})>C\tilde{C}\frac{2^{m+2}}{\tilde{\alpha}}\,\cdot$ (247)
By the weak convergence (232) we know that $\delta_{n,N}\rightarrow 0$, and so we can choose $n$ big enough to guarantee
$\delta_{n,N}<\frac{\tilde{\alpha}}{6L_{N}}\,\cdot$ (248)
With these choices we eventually get
$\tilde{E}_{r_{n}}(f_{k,r_{n}}-f)>C,$ (249)
which is a contradiction.
Step 2. Now we show that $\tilde{\lambda}_{k,r_{n}}\rightarrow\mu_{k}$, where $\mu_{k}$ is the $k$-th Neumann eigenvalue. Let $r_{n}\rightarrow 0$. We know by Proposition 6.2 that there exist eigenfunctions $f_{0},...,f_{k}$ with Neumann eigenvalue $\lambda_{0},...,\lambda_{k}$ such that
$f_{i,r_{n}}\stackrel{{\scriptstyle L^{2}}}{{\rightarrow}}f_{i},\quad\quad\forall i\in\{0,...,k\},$ (250)
$\tilde{\lambda}_{k,r_{n}}\rightarrow C_{m}\lambda_{k},$ (251)
and
$\lambda_{i}\leq\lambda_{k}\quad\quad\forall i\in\{0,...,k\}.$ (252)
Since $\langle f_{i,r_{n}},f_{j,r_{n}}\rangle=\delta_{i,j}$, we also have by strong convergence that $\langle f_{i},f_{j}\rangle=\delta_{i,j}$. Thus we have that
$V_{k+1}\coloneqq\mathrm{Span}(f_{0},\ldots,f_{k})\in\mathcal{G}_{k+1}(L^{2}(M,\mathrm{vol}_{g})),$ (253)
and so by equation (252), we conclude
$C_{m}\mu_{k}\leq\max_{f\in V_{k+1}}\frac{\langle\nabla f,\nabla f\rangle}{\|f\|_{L^{2}}}=C_{m}\lambda_{k}=\lim_{n}\tilde{\lambda}_{k,r_{n}}.$ (254)
This shows that $\liminf_{r\rightarrow 0}\tilde{\lambda}_{k,r_{n}}\geq C_{m}\mu_{k}$
To prove $\limsup_{r\rightarrow 0}\tilde{\lambda}_{k,r}\leq C_{m}\mu_{k}$, let $\{f_{0},\ldots,f_{k}\}$ be an $\langle\cdot,\cdot\rangle_{2}$-orthonormal family of Laplace (resp. Neumann) eigenfunctions associated with the eigenvalues $\{\mu_{0},\ldots,\mu_{k}\}$ respectively satisfying $\mu_{0}\leq...\leq\mu_{k}$. By elliptic regularity, we know that these functions belong to $C^{\infty}(M)$. Then Proposition 5.1 implies that given $\epsilon>0$, there exists $r_{\epsilon}>0$ such that for $r\in(0,r_{\epsilon})$,
$\left|\langle-\tilde{\Delta}_{r}f_{i},f_{j}\rangle_{2}-\delta_{i,j}C_{m}\mu_{j}\right|<\epsilon$ (255)
where $\delta_{i,j}$ is the usual Kronecker delta. Set $U\coloneqq\mathrm{Span}\left(f_{0},...,f_{k}\right)$ and
$v\coloneqq\sum_{i=1}^{k}a_{i}\psi_{i}$ (256)
for some $a=(a_{1},\ldots,a_{k})\in\mathbb{S}^{k-1}$. Then
$\left|\langle-\tilde{\Delta}_{r}v,v\rangle-\sum_{i=1}^{k}a_{i}^{2}C_{m}\mu_{i}\right|=\left|\sum_{i,j=1}^{k}a_{i}a_{j}\langle-\tilde{\Delta}_{r}f_{i},f_{j}\rangle-\sum_{i=1}^{k}a_{i}^{2}C_{m}\mu_{i}\right|\leq k^{2}\epsilon.$ (257)
Since $U$ is a $k+1$-dimensional subspace, we conclude that
$\tilde{\lambda}_{k,r}\leq\max_{v\in U}\frac{\langle-\tilde{\Delta}_{r}v,v\rangle}{\|v\|_{2}^{2}}\leq\max_{a\in\mathbb{S}^{k}}\sum_{i=1}^{k}a_{i}^{2}\mu_{i}+k^{2}\epsilon\leq\mu_{k}+k^{2}\epsilon.$ (258)
Take the limit superior as $r\to 0$ and then let $\epsilon\to 0$ to obtain $\limsup_{r\rightarrow 0}\tilde{\lambda}_{k,r}\leq C_{m}\mu_{k}$. Combined with Corollary 3.8, the latter implies the existence of $r_{k}>0$ such that $\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\big{)}\geq\mu_{k}+1\geq\tilde{\lambda}_{k,r}$ for any $r\in(0,r_{k})$, so that $\tilde{\lambda}_{k,r}$ indeed coincides with $\lambda_{k}(-\tilde{\Delta}_{r})$. ∎

### v3

@@
 These form a non-decreasing sequence of non-negative numbers.
-Our first main result states that these numbers eventually correspond to isolated eigenvalues of $-\tilde{\Delta}_{r}$ when $(X,\mathsf{d},\mu)$ is compact and uniformly locally doubling (Definition 2.1).
+Our first main result states that these numbers eventually correspond to isolated eigenvalues of $\tilde{\Delta}_{r}$ when $(X,\mathsf{d},\mu)$ is compact and uniformly locally doubling (Definition 2.8).
 Theorem 1.
 Let $(X,\mathsf{d},\mu)$ be a compact uniformly locally doubling metric measure space.
-For any integer $k\geq 2$, there exists $r_{k}>0$ such that for any $r\in(0,r_{k})$, the operator $-\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues
+For any integer $k\geq 2$, there exists $r_{k}>0$ such that for any $r\in(0,r_{k})$, the operator $\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues
 $0=\lambda_{0}(-\tilde{\Delta}_{r})<\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$
@@
 After the previous $L^{2}$-convergence result, we address the question of spectral convergence, that is to say, the convergence of the associated eigenvalues and eigenfunctions.
-In this regard we show that, for any $k\in\mathbb{N}$, the function $r\mapsto\tilde{\lambda}_{k,r}$ is bounded in a neighborhood of 00, as proved in the course of Theorem 1.
-This ensures that the $k$-th lowest eigenvalue of the operator $-\tilde{\Delta}_{r}$, which we denote $\lambda_{k}(-\tilde{\Delta}_{r})$, exists for small enough $r$, and that it coincides with $\tilde{\lambda}_{k,r}$.
-Let $f_{k,r}$ be an $L^{2}$-normalized eigenfunction of $-\tilde{\Delta}_{r}$ associated with $\lambda_{k}(-\tilde{\Delta}_{r})$.
+In this regard we show that, for any $k\in\mathbb{N}$, the function $r\mapsto\tilde{\lambda}_{k,r}$ is bounded in a neighborhood of 0, as proved in the course of Theorem 1.
+This ensures that the $k$-th lowest eigenvalue of the operator $\tilde{\Delta}_{r}$, which we denote $\lambda_{k}(-\tilde{\Delta}_{r})$, exists for small enough $r$, and that it coincides with $\tilde{\lambda}_{k,r}$.
+Let $f_{k,r}$ be an $L^{2}$-normalized eigenfunction of $\tilde{\Delta}_{r}$ associated with $\lambda_{k}(-\tilde{\Delta}_{r})$.
 Recall that if $\partial M=\emptyset$ (resp.
@@
 Neumann) eigenvalue of $(M,g)$ is a number $\mu\geq 0$ for which there exists an associated eigenfuction $f\in\mathcal{C}^{\infty}(M)$ (resp.
-$\mathcal{C}_{\nu}^{\infty}(M)$) of $-\Delta_{g}$, i.e.
-$-\mu f=\Delta_{g}f$.
+$\mathcal{C}_{\nu}^{\infty}(M)$) of $\Delta_{g}$, i.e.
+$\mu f=\Delta_{g}f$.
 Theorem 3.
@@
 G0DBZ23N.
-They thank the anonymous reviewers for their constructive suggestions, and Dmitri Pavlov for clarifying a point related to Lemma 2.1.
+They thank the anonymous reviewers for their numerous and constructive suggestions, and Dmitri Pavlov for clarifying a point related to Lemma 2.1.
 2. Averaging-like operators
@@
 $\left.\mu\right|_{B_{N}(o)}(\cup_{n}B_{\epsilon}(y_{n}))=\alpha_{\epsilon,N}.$ (11)
-Now we prove $\alpha_{\epsilon,N}=\mu(B_{N}(z))<\infty$.
-If $\alpha_{\epsilon,N}<\mu(B_{N}(z))$, then $\mu(B_{N}(z)\backslash\cup_{n}B_{\epsilon}(y_{n}))>0$, where $\{y_{n}\}_{n\in\mathbb{N}}$ is a maximizer of (8).
-Since the measure is inner regular, there must exist some compact $K\subset B_{N}(z)\backslash\cup_{n}B_{\epsilon}(y_{n})$ such that
+[…]
 $\mu(K)>0.$ (12)
@@
 However, there may be no uniform $r>0$ for which the integral $A_{r}u(x)$ is well-defined for every $x\in X$.
-In the next lemma, we provide a mild condition under which $A_{r}$ defines a bounded linear operator between Lebesgue spaces.
+Let us also set
+$a^{*}_{r}(x,y)\coloneqq a_{r}(y,x)=\frac{1_{B_{r}(x)}(y)}{V(y,r)}$
+for any $x,y\in X$ and $r>0$.
+Consider $u\in L^{0}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$.
+For any $x\in X$ and $r>0$ such that $v$ is $\mu$-integrable on $B_{r}(x)$, set
+$A^{*}_{r}u(x)\coloneqq\int_{B_{r}(x)}\frac{u(y)\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}=\int_{X}a_{r}^{*}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$
+Notice that, just like $A_{r}u(x)$, $A^{*}_{r}u(x)$ may not make sense uniformly with respect to $x\in X$.
+For any $r>0$, we introduce the following conditions:
+$\|A_{r}^{*}1\|_{\infty}<+\infty,$ ($\mathrm{I}_{r}$)
+$V(\cdot,r)^{-1}\in L^{1}(X,\mu).$ ($\mathrm{II}_{r}$)
+Note that ($\mathrm{II}_{r}$) implies ($\mathrm{I}_{r}$) since
+$\|A_{r}^{*}1\|_{\infty}=\sup_{x\in X}|A_{r}^{*}1(x)|=\sup_{x\in X}\int_{B_{r}(x)}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\leq\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\,\cdot$
+In the next lemma, we discuss the boundedness and the compactness of the averaging operator $A_{r}$ acting on Lebesgue spaces.
 Lemma 2.2.
-Assume that there exists $r>0$ such that
-$V(\cdot,r)^{-1}\in L^{1}(X,\mu).$ ($\mathrm{A}_{r}$)
-Then for any $p\in[1,+\infty]$ the linear operator $A_{r}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
-$\|A_{r}\|_{p\to p}\leq\|V(\cdot,r)^{-1}\|_{1}^{1/p}\cdot$
-Moreover, $A_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
+Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then for any $p\in[1,+\infty]$ the linear operator $A_{r}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
+$\|A_{r}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{1/p}\cdot$
+Moreover, if ($\mathrm{II}_{r}$) holds, then $A_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
 Proof.
-The case $p=+\infty$ is obvious, hence we assume $p<+\infty$. Let $u\in L^{p}(X,\mu)$. By Jensen’s inequality, for any $x\in X$,
+The case $p=+\infty$ is obvious and holds regardless of ($\mathrm{I}_{r}$). Let us assume that $p<+\infty$. Let $u\in L^{p}(X,\mu)$. By Jensen’s inequality, for any $x\in X$,
 $|A_{r}u(x)|^{p}\leq\left(\fint_{B_{r}(x)}|u|^{p}\mathop{}\!\mathrm{d}\mu\right).$
@@
 $\|A_{r}u\|_{p}^{p}$ $\leq\int_{X}\frac{1}{V(x,r)}\int_{B_{r}(x)}|u(y)|^{p}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
-$\leq\left(\int_{X}\frac{1}{V(x,r)}\mathop{}\!\mathrm{d}\mu(x)\right)\|u\|_{p}^{p}.$
-Since
+$=\int_{X}\int_{X}\frac{1}{V(x,r)}\underbrace{1_{B_{r}(x)}(y)}_{=1_{B_{r}(y)}(x)}|u(y)|^{p}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
+$=\int_{X}|u(y)|^{p}\underbrace{\int_{B_{r}(y)}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}}_{=A_{r}^{*}1(y)}\mathop{}\!\mathrm{d}\mu(y)\leq\|A_{r}^{*}1\|_{\infty}\|u\|_{p}^{p}$
+where we have used the Fubini–Tonelli theorem to get the second equality and ($\mathrm{I}_{r}$) for the last inequality.
+Let us now assume that ($\mathrm{II}_{r}$) holds. Since
 $\int_{X}\int_{X}a_{r}^{2}(x,y)\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)=\int_{X}\frac{1}{V(x,r)}\fint_{B_{r}(x)}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)=\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}$
-we obtain that $A_{r}$ is a Hilbert-Schmidt integral operator acting on the separable space $L^{2}(X,\mu)$ (recall Lemma 2.1); in particular, it is compact [RS80, Section IV.6]. ∎
-Lemma 2.3.
-Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{A}_{r}$) holds for any $r>0$.
-Proof.
-Consider $r>0$ and a finite cover $\{B_{r/2}(x_{i})\}$ of $X$.
-For any $x\in X$, there exists $i$ such that $x\in B_{r/2}(x_{i})$.
-Then $B_{r}(x)$ contains $B_{r/2}(x_{i})$ so $V(x,r)\geq V(x_{i},r/2)\geq\min_{j}V(x_{j},r/2)>0$.
-Thus
-$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{\min_{j}V(x_{j},r/2)}<+\infty.$
-Remark 2.4.
-Notice that if $\mu(X)=+\infty$ and $M(r)<+\infty$ then ($\mathrm{A}_{r}$) cannot hold :
-$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\geq\frac{\mu(X)}{M(r)}=+\infty.$
-This happens, for instance, on $\mathbb{R}^{n}$ endowed with the Euclidean distance and the Lebesgue measure.
-More generally, this property holds on any locally compact topological group endowed with a left-invariant metric compatible with the Haar measure and with infinite volume (see [Str74, Lemma 1] for more details about these spaces).
-Moreover, if $\mu(X)<\infty$ and $m(r)>0$, then ($\mathrm{A}_{r}$) always holds :
-$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{m(r)}<+\infty.$
-In this regard, observe that if $X$ is not totally bounded, then there exist $r>0$ small enough and a countable family of disjoint balls $\{B_{r}(x_{i})\}$ in $X$, so that $\mu(X)<\infty$ and $m(r)>0$ cannot hold simultaneously:
-$\mu(X)\geq\sum_{i}V(x_{i},r).$
+we obtain that $A_{r}$ is a Hilbert-Schmidt integral operator acting on the separable space $L^{2}(X,\mu)$ (recall Lemma 2.1); in particular, $A_{r}$ is compact [RS80, Section IV.6]. ∎
 In the next statement, we provide an alternative way to prove the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to itself.
-This goes through the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to the space of continuous functions $\mathcal{C}(X)$ which we obtain for $X$ compact and satisfying the following condition:
+This goes through the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to the space of continuous functions $\mathcal{C}(X)$ which we obtain for compact spaces $X$ satisfying the following condition:
 $\sup_{x\in X}\mu(S_{r}(x))=0$ ($\mathrm{S}_{r}$)
@@
 Proof.
-[…]
-[…]
+We start by noticing that if $u\in L^{2}(X,\mu)$, then $A_{r}(u)$ is continuous.
+This follows from $V(\cdot,r)^{-1}$ and $\int_{B_{r}(x)}u(y)d\mu(y)$ being continuous.
+The former holds by assumption.
+To prove the latter, assume that $\|u\|_{L^{2}(X)}=1$.
+Then for any $x,z\in X$,
+$\left|\int_{B^{r}(x)}u(y)d\mu(y)-\int_{B^{r}(z)}u(y)d\mu(y)\right|$ $\leq\left\|1_{B_{r}(x)}-1_{B_{r}(z)}\right\|_{L^{2}(X)}\left\|u\right\|_{L^{2}(X)}$ (21)
 $\leq\mu(B_{r+d(x,z)}(x)-B_{r-d(x,z)}(x))^{\frac{1}{2}},$ (22)
@@
 2.2. Adjoint
-For any $x,y\in X$ and $r>0$, set
-$a^{*}_{r}(x,y)\coloneqq a_{r}(y,x)=\frac{1_{B_{r}(x)}(y)}{V(y,r)}\,\cdot$
-Consider $u\in L^{0}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $v$ is $\mu$-integrable on $B_{r}(x)$, set
-$A^{*}_{r}u(x)\coloneqq\int_{B_{r}(x)}\frac{u(y)\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}=\int_{X}a_{r}^{*}(x,y)\mathop{}\!\mathrm{d}\mu(y).$
-Notice that, just like $A_{r}u(x)$ in the previous subsection, $A^{*}_{r}u(x)$ may not make sense uniformly with respect to $x\in X$.
-Let us begin with a simple observation.
+Let us focus now on the boundedness and the compactness of the adjoint operator $A_{r}^{*}$. We begin with a simple observation.
 Lemma 2.4.
-The operator $A_{r}^{*}:L^{1}(X,\mu)\to L^{1}(X,\mu)$ is a contraction.
+The operator $A_{r}^{*}:L^{1}(X,\mu)\to L^{1}(X,\mu)$ is a contraction for any $r>0$.
 Proof.
@@
 where we used the Fubini–Tonelli theorem to get the penultimate equality. ∎
-The next lemma gives a mild sufficient condition ensuring the well-posedeness and boundedness of the linear operator $A_{r}^{*}$ between $L^{p}$ spaces for $p>1$.
+We continue with the next lemma which covers the case $p>1$.
 Lemma 2.5.
-Assume that there exists $r>0$ such that
-$\left.\begin{array}[]{l}V(\cdot,r)^{-1}\in L^{\infty}(B_{r}(x),\mu)\qquad\forall x\in X,\\ \|A_{r}^{*}1\|_{\infty}<+\infty.\end{array}\right\}$ ($\mathrm{A}_{r}^{*}$)
-Then for any $p\in[1,+\infty]$, the linear operator $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
-$\|A_{r}^{*}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{(p-1)/p}.$
-Moreover, if ($\mathrm{A}_{r}$) additionally holds, then $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is the adjoint of $A_{r}:L^{q}(X,\mu)\to L^{q}(X,\mu)$ for $q\in[1,+\infty]$ such that $11/p+1/q=$ 1, and the operator $A_{r}^{*}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
+Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then for any $p\in[1,+\infty]$, the linear operator $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
+$\|A_{r}^{*}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{(p-1)/p}$
+Moreover, this operator is the adjoint of $A_{r}:L^{q}(X,\mu)\to L^{q}(X,\mu)$ for $q\in[1,+\infty]$ such that $1/p+1/q=1$.
+Lastly, if ($\mathrm{II}_{r}$) holds, then the operator $A_{r}^{*}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
 Proof.
-For the proof of the first assertion, consider $u\in L^{\infty}(X,\mu)$. Then for any $x\in X$,
+For the proof of the first assertion, consider $u\in L^{\infty}(X,\mu)$. Thanks to ($\mathrm{I}_{r}$), for $\mu$-a.e. $x\in X$,
 $|A_{r}^{*}u|(x)$ $\leq\int_{B_{r}(x)}\frac{|u(y)|}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq\|u\|_{\infty}\int_{B_{r}(x)}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\leq\|A_{r}^{*}1\|_{\infty}\|u\|_{\infty}.$
@@
 The conclusion for $p\in(1,+\infty)$ follows from the Riesz-Thorin theorem and Lemma 2.4.
-Assume now that ($\mathrm{A}_{r}$) holds, and consider $u\in L^{p}(X,\mu)$ and $v\in L^{q}(X,\mu)$. Then
+Let us prove that $A_{r}$ and $A_{r}^{*}$ are adjoint of each other. Consider $u\in L^{p}(X,\mu)$ and $v\in L^{q}(X,\mu)$. Then
 $\int_{X}A_{r}^{*}u(x)\,v(x)\mathop{}\!\mathrm{d}\mu(x)$ $=\int_{X}\int_{X}1_{B_{r}(x)}(y)\,\frac{u(y)}{V(y,r)}\,v(x)\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
@@
 $=\int_{X}u(y)A_{r}v(y)\mathop{}\!\mathrm{d}\mu(y)$
-where we used the Fubini–Tonelli to get the second equality, and the symmetry of the distance function through the equality $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ to get the last one.
-As for the compactness of $A_{r}^{*}$, the result is a direct consequence of the Schauder theorem for compact operators which can be applied thanks to Lemma 2.2.
+where we used the Fubini–Tonelli to get the second equality, and the equality $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ to get the last one.
+As for the compactness of $A_{r}^{*}$ under ($\mathrm{II}_{r}$), this result is a direct consequence of the Schauder theorem for compact operators which can be applied thanks to Lemma 2.2.
 ∎
+2.3. Discussion on the assumptions
+Let us discuss the validity of ($\mathrm{I}_{r}$) and ($\mathrm{II}_{r}$).
+Recall first that : ($\mathrm{II}_{r}$) $\Rightarrow$ ($\mathrm{I}_{r}$).
+Both properties hold on totally bounded spaces, as seen in the next lemma.
 Lemma 2.6.
-Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{A}_{r}^{*}$) holds for any $r>0$.
+Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{II}_{r}$) (and then ($\mathrm{I}_{r}$)) holds for any $r>0$.
 Proof.
-As seen in the proof of Lemma 2.3, total boundedness implies $m(r)>0$ for any $r>0$ hence $V(\cdot,r)^{-1}\in L^{\infty}(B_{r}(x),\mu)$ for any $x\in X$.
-Total boundedness also implies that $\mu(X)<+\infty$, thus $\|A_{r}^{*}1\|_{\infty}\leq\mu(X)/m(r)<+\infty$.
-∎
+Consider $r>0$ and a finite cover $\{B_{r/2}(x_{i})\}$ of $X$.
+For any $x\in X$, there exists $i$ such that $x\in B_{r/2}(x_{i})$.
+Then $B_{r}(x)$ contains $B_{r/2}(x_{i})$ so $V(x,r)\geq V(x_{i},r/2)\geq\min_{j}V(x_{j},r/2)>0$.
+Thus
+$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{\min_{j}V(x_{j},r/2)}<+\infty.$
 Remark 2.7.
-Note that if $m(r)>0$ and $M(r)<+\infty$, then ($\mathrm{A}_{r}^{*}$) always holds with
+If $\mu(X)=+\infty$ and $M(r)<+\infty$ then ($\mathrm{II}_{r}$) cannot hold :
+$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\geq\frac{\mu(X)}{M(r)}=+\infty.$
+This happens, for instance, on $\mathbb{R}^{n}$ endowed with the Euclidean distance and the Lebesgue measure.
+More generally, this property cannot hold on a locally compact topological group endowed with a left-invariant metric compatible with the Haar measure and with infinite volume (see [Str74, Lemma 1] for more details about these spaces).
+If $\mu(X)<\infty$ and $m(r)>0$, then ($\mathrm{II}_{r}$) always holds :
+$\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}\leq\frac{\mu(X)}{m(r)}<+\infty.$
+In this regard, observe that if $X$ is not totally bounded, then there exist $r>0$ small enough and a countable family of disjoint balls $\{B_{r}(x_{i})\}$ in $X$, so that $\mu(X)<\infty$ and $m(r)>0$ cannot hold simultaneously:
+$\mu(X)\geq\sum_{i}V(x_{i},r).$
+Let us now focus on ($\mathrm{I}_{r}$).
+We show below that this condition holds on so-called doubling spaces.
+Let us recall this classical property and its uniform local variant, see e.g.
+[HKST15] for more details.
+Definition 2.8.
+The space $(X,\mathsf{d},\mu)$ is called globally doubling if there exists $C>0$ such that for any $x\in X$ and $r>0$,
+$V(x,2r)\leq CV(x,r).$ (23)
+It is called uniformly locally doubling if there exist $C,r_{0}>0$ such that (23) holds for any $x\in X$ and $r\in(0,r_{0})$.
+The celebrated Bishop–Gromov theorem (see e.g.
+[Cha95, Theorem III.4.5]) implies that any complete Riemannian manifold with a uniform lower bound on the Ricci curvature is uniformly uniformly locally doubling, and globally doubling if the uniform bound is non-negative.
+This is also true for metric spaces with generalized sectional curvature bounded from below in the sense of Alexandrov [BBI01, Theorem 10.6.6] and $\mathrm{CD}(K,N)$ metric measure space [Vil09, Corollary 30.14].
+The next lemma relates the uniformly local doubling condition with ($\mathrm{I}_{r}$).
+Lemma 2.9.
+Let $(X,\mathsf{d},\mu)$ be uniformly locally doubling with parameters $C,r_{0}$. Then ($\mathrm{I}_{r}$) holds with $\|A_{r}^{*}1\|_{\infty}\leq C$ for any $r\in(0,r_{0})$.
+Proof.
+For any $x\in X$ and $r\in(0,r_{0})$, the triangle inequality yields that $B_{r}(x)\subset B_{2r}(y)$ for any $y\in B_{r}(x)$. Then
+$A_{r}^{*}1(x)=\fint_{B_{r}(x)}\frac{V(x,r)}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq\fint_{B_{r}(x)}\frac{V(y,2r)}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq C.$
+Remark 2.10.
+Of course if $(X,\mathsf{d},\mu)$ is globally doubling with constant $C$, then ($\mathrm{I}_{r}$) holds with $\|A_{r}^{*}1\|_{\infty}\leq C$ for any $r>0$.
+Remark 2.11.
+The previous result notably implies that ($\mathrm{I}_{r}$) may hold in situations where ($\mathrm{II}_{r}$) does not.
+This happens e.g.
+on a non-compact Riemannian manifold $(M,g)$ with non-negative Ricci curvature endowed with its canonical Riemannian distance $\mathsf{d}$ and volume measure $\mu$.
+Indeed, such a space has infinite volume, and the Bishop–Gromov theorem implies that $M(r)\leq\mathbb{V}^{n}(r)$ for any $r>0$, where $\mathbb{V}^{n}(r)$ is the Lebesgue measure of an Euclidean ball of radius $r$ in $\mathbb{R}^{n}$.
+From Remark 2.7, we get that $M$ cannot satisfy ($\mathrm{II}_{r}$) for any $r>0$, while it does satisfies ($\mathrm{I}_{r}$) thanks to the global doubling condition.
+We conclude this discussion with two final remarks. First, if $m(r)>0$ and $M(r)<+\infty$, then ($\mathrm{I}_{r}$) always holds with
 $\|A_{r}^{*}1\|_{\infty}\leq\frac{M(r)}{m(r)}\,\cdot$
-This notably happens on locally compact topological groups endowed with a left-invariant distance and their Haar measure, compare with Remark 2.4.
-Furthermore, observe that $A_{r}^{*}1\in L^{\infty}(X,\mu)$ can be seen as a weak variant of the comparability conditions introduced in [Ald18, MT23].
+This happens on locally compact topological groups endowed with a left-invariant distance and their Haar measure, compare with Remark 2.7.
+Secondly, ($\mathrm{I}_{r}$) can easily be seen as a weak variant of the comparability conditions introduced in [Ald18, MT23].
 2.4. Symmetrization
@@
 Corollary 2.12.
-Assume that there exists $r>0$ such that ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) hold. Then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is a compact self-adjoint operator such that
-$\|\tilde{A}_{r}\|_{2\to 2}\leq\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}$
-2.4. Doubling condition
-Let us recall the classical doubling condition and its uniform local variant, see e.g. [HKST15] for more details.
-Definition 2.1.
-The space $(X,\mathsf{d},\mu)$ is called doubling if there exists $C>0$ such that for any $x\in X$ and $r>0$,
-$V(x,2r)\leq CV(x,r).$ (24)
-It is called uniformly uniformly locally doubling if there exist $C,r_{0}>0$ such that (24) holds for any $x\in X$ and $r\in(0,r_{0})$.
-It is worth pointing out that the celebrated Bishop–Gromov theorem, see e.g.
-[Cha95, Theorem III.4.5], implies that any complete Riemannian manifold with a uniform lower bound on the Ricci curvature is uniformly uniformly locally doubling.
-This is also true for metric spaces with generalized sectional curvature bounded from below in the sense of Alexandrov [BBI01, Theorem 10.6.6] and $\mathrm{CD}(K,N)$ metric measure space [Vil09, Corollary 30.14].
+Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is a self-adjoint operator such that
+$\|\tilde{A}_{r}\|_{2\to 2}\leq\|A_{r}^{*}1\|_{\infty}^{1/2}.$
+Moreover, if ($\mathrm{II}_{r}$) holds, then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
 3. Symmetrized AMV operators
@@
 3.1. Definitions
-For this subsection, we consider a metric measure space $(X,\mathsf{d},\mu)$ satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$.
+For this subsection, we consider a metric measure space $(X,\mathsf{d},\mu)$ satisfying ($\mathrm{I}_{r}$) for some fixed $r>0$.
 Definition 3.1.
@@
 $\tilde{\Delta}_{r}$ is a bounded, self-adjoint operator acting on $L^{2}(X,\mu)$ with
-$\|\tilde{\Delta}_{r}\|_{2\to 2}\leq\frac{1}{r^{2}}\left(\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}+c(r)+1\right).$ (25)
+$\|\tilde{\Delta}_{r}\|_{2\to 2}\leq\frac{1}{2r^{2}}\left(2\|A_{r}^{*}1\|_{\infty}^{1/2}+\|A_{r}^{*}1\|_{\infty}+1\right).$ (25)
 Proof.
-The self-adjointness of $\tilde{\Delta}_{r}$ is obvious because $\tilde{A}_{r}$ and $[\tilde{A}_{r}1]\mathrm{I}$ are self-adjoint too. The boundedness is a consequence of Lemma 2.10. Indeed,
+The self-adjointness of $\tilde{\Delta}_{r}$ is obvious because $\tilde{A}_{r}$ and $\tilde{A}_{r}1]\mathrm{I}$ are self-adjoint too. The boundedness is a consequence of Lemma 2.12. Indeed,
 $\|\tilde{\Delta}_{r}\|_{2\to 2}$ $\leq\frac{1}{r^{2}}\left(\|\tilde{A}_{r}\|_{2\to 2}+\|[\tilde{A}_{r}1]I\|_{2\to 2}\right)$
-$\leq\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\|\tilde{A}_{r}1\|_{\infty}\underbrace{\|I\|_{2\to 2}}_{=1}\right)$
-$\leq\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\frac{\|A_{r}1\|_{\infty}+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
-$=\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\frac{1+c(r)}{2}\right)$
+$\leq\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\|\tilde{A}_{r}1\|_{\infty}\underbrace{\|I\|_{2\to 2}}_{=1}\right)$
+$\leq\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\frac{\|A_{r}1\|_{\infty}+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
+$=\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\frac{1+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
 hence $\tilde{\Delta}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is bounded and (25) holds. ∎
@@
 $\tilde{E}_{r}(f,\psi)$ $\coloneqq\frac{1}{4}\int_{X}\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)$
-×$\qquad\qquad\qquad\qquad\times\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
+$\qquad\qquad\qquad\qquad\times\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
+Remark 3.5.
+A suitable use of the Fubini–Tonelli theorem shows that the energy functional $\tilde{E}_{r}(f)$ equals the approximate Korevaar–Schoen energy [KS93]
+$\frac{1}{2}\int_{X}\fint_{B_{r}(x)}\frac{|f(y)-f(x)|^{2}}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
 Remark 3.6.
@@
 Remark 3.8.
-Observe that (26) implies that $-\tilde{\Delta}_{r}$ is a non-negative operator, since for any $f\in L^{2}(X,\mu)$,
+Observe that (26) implies that $\tilde{\Delta}_{r}$ is a non-negative operator, since for any $f\in L^{2}(X,\mu)$,
 $\langle-\tilde{\Delta}_{r}f,f\rangle_{L^{2}}=\tilde{E}_{r}(f)\geq 0$
@@
 Definition 3.9.
-We let $\sigma(-\tilde{\Delta}_{r})$ denote the spectrum of $-\tilde{\Delta}_{r}$, that is to say, the set of elements $\lambda\in\mathbb{C}^{*}$ such that $-\tilde{\Delta}_{r}-\lambda\mathrm{I}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is not a bijection.
+We let $\sigma(-\tilde{\Delta}_{r})$ denote the spectrum of $\tilde{\Delta}_{r}$, that is to say, the set of elements $\lambda\in\mathbb{C}^{*}$ such that $\tilde{\Delta}_{r}-\lambda\mathrm{I}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is not a bijection.
 It is well-known from classical functional analysis that the spectrum $\sigma(T)$ of a bounded operator $T$ acting on a Banach space $E$ can be decomposed as
@@
   • $\sigma_{\text{p}}(T)$ is the point spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ such that $(T-\lambda\mathrm{I})f=0$ for some non-zero $f\in E$, in which case $\lambda$ is called an eigenvalue and $f$ an eigenvector of $T$,
-  • $\sigma_{\color[rgb]{.75,0,.25}\text{c}}(T)$ is the compression spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ whose conjugate $\bar{\lambda}$ is an eigenvalue of the adjoint $T^{*}$,
+  • $\sigma_{\text{c}}(T)$ is the compression spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ whose conjugate $\bar{\lambda}$ is an eigenvalue of the adjoint $T^{*}$,
   • $\sigma_{\text{a}}(T)$ is the approximate point spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ for which there exists $(f_{n})\subset E$ with $\|f_{n}\|=1$ for any $n$ such that $\|(T-\lambda\mathrm{I})f_{n}\|\to 0$.
-Since $-\tilde{\Delta}_{r}$ is self-adjoint and non-negative, we know that
+Since $\tilde{\Delta}_{r}$ is self-adjoint and non-negative, we know that
 $\sigma(-\tilde{\Delta}_{r})\subset[0,+\infty].$
@@
 Remark 3.11.
-Let $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ denote the essential spectrum of $-\tilde{\Delta}_{r}$, i.e.
-the closed subset of $\sigma(-\tilde{\Delta}_{r})$ made of those $\lambda$ such that $-\tilde{\Delta}_{r}-\lambda\mathrm{I}$ is not a Fredholm operator.
-Since $-\tilde{\Delta}_{r}$ is self-adjoint, the Fischer–Polyà minimum-maximum principle (see e.g.
-[WS72, p.12]) asserts that if there exists a positive integer $N$ such that $\tilde{\lambda}_{N,r}<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ then $-\tilde{\Delta}_{r}$ admits $N+1$ isolated eigenvalues
+Let $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ denote the essential spectrum of $\tilde{\Delta}_{r}$, i.e.
+the closed subset of $\sigma(-\tilde{\Delta}_{r})$ made of those $\lambda$ such that $\tilde{\Delta}_{r}-\lambda\mathrm{I}$ is not a Fredholm operator.
+Since $\tilde{\Delta}_{r}$ is self-adjoint, the Fischer–Polyà minimum-maximum principle (see e.g.
+[WS72, p.12]) asserts that if there exists a positive integer $N$ such that $\tilde{\lambda}_{N,r}<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ then $\tilde{\Delta}_{r}$ admits $N+1$ isolated eigenvalues
 $\lambda_{0}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{N}(-\tilde{\Delta}_{r})<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$
@@
 3.2. Spectral properties
-Our first spectral result on $\tilde{\Delta}_{r}$ is the following.
+Our first spectral result on $\tilde{\Delta}_{r}$ is the following. Note that we need the compactness of $\tilde{A}_{r}$ here, thus we assume ($\mathrm{II}_{r}$).
 Proposition 3.12.
-Let $(X,\mathsf{d},\mu)$ be a metric measure space satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$. Assume that $\lambda\in\sigma(-\tilde{\Delta}_{r})$ satisfies
+Let $(X,\mathsf{d},\mu)$ be a metric measure space satisfying ($\mathrm{II}_{r}$) for some fixed $r>0$. Assume that $\lambda\in\sigma(-\tilde{\Delta}_{r})$ satisfies
 $\lambda<\inf_{y\in X}[\tilde{A}_{r}1](y)/{r^{2}}.$ (29)
-Then $\lambda$ is an isolated eigenvalue of $-\tilde{\Delta}_{r}$ which does not belong to $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$.
+Then $\lambda$ is an isolated eigenvalue of $\tilde{\Delta}_{r}$ which does not belong to $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$.
 Proof.
@@
 $\delta\|f_{n}-\frac{g}{b_{r}}\|_{L^{2}(X)}\leq\|b_{r}f_{n}-g\|_{L^{2}(X)}\rightarrow 0.$ (32)
-Thus $f_{n}$ converges in $L^{2}(X,\mu)$ to the limit function $f$. Using continuity of $-\tilde{\Delta}_{r}$ we conclude that
-$-\tilde{\Delta}_{r}f=\lambda f,$ (33)
+Thus $f_{n}$ converges in $L^{2}(X,\mu)$ to the limit function $f$. Using continuity of $\tilde{\Delta}_{r}$ we conclude that
+$\tilde{\Delta}_{r}f=\lambda f,$ (33)
 and thus $\lambda$ is in the point spectrum.
@@
 From the previous paragraph, we know that $\lambda_{n}$ is in the point spectrum, thus there exists $f_{n}\in L^{2}(X,\mu)$ satisfying $\|f_{n}\|_{L^{2}(X)}=1$, such that
-$-\tilde{\Delta}_{r}f_{n}=\lambda_{n}f_{n}.$ (35)
+$\tilde{\Delta}_{r}f_{n}=\lambda_{n}f_{n}.$ (35)
 This can be written as
-$-\frac{\tilde{A}_{r}}{r^{2}}f_{n}=(\lambda_{n}-\frac{[\tilde{A}_{r}1]}{r^{2}})f_{n}.$ (36)
+$\frac{\tilde{A}_{r}}{r^{2}}f_{n}=(\lambda_{n}-\frac{[\tilde{A}_{r}1]}{r^{2}})f_{n}.$ (36)
 Using compactness of $\tilde{A}_{r}$, we know that $\tilde{A}_{r}f_{n}$ converges up to a subsequence.
@@
 This shows that $\lambda$ is an isolated point of $\sigma(-\tilde{\Delta}_{r})$ finishing the first part of the proof.
-Let us now prove that $-\tilde{\Delta}_{r}-\lambda I$ is a Fredholm operator.
+Let us now prove that $\tilde{\Delta}_{r}-\lambda I$ is a Fredholm operator.
 To show that $\ker(-\tilde{\Delta}_{r}-\lambda I)$ is finite dimensional we proceed by contradiction.
@@
 Thus
-$-\frac{\tilde{A}_{r}f_{n}}{r^{2}}=\left(\lambda-\frac{[\tilde{A}_{r}1]}{r^{2}}\right)f_{n}.$ (40)
+$\frac{\tilde{A}_{r}f_{n}}{r^{2}}=\left(\lambda-\frac{[\tilde{A}_{r}1]}{r^{2}}\right)f_{n}.$ (40)
 Similar to before we can use compactness of $\tilde{A}_{r}$ and the condition on $\lambda$ to conclude that $f_{n}$ converges in $L^{2}(X,\mu)$ up to a subsequence.
 However, this is prevented by $\langle f_{n},f_{j}\rangle=\delta_{n,j}$.
-To show that the image of $-\tilde{\Delta}_{r}-\lambda I$ is closed, consider a sequence $g_{n}:=(-\tilde{\Delta}_{r}-\lambda I)(f_{n})$ such that $g_{n}\rightarrow g$.
+To show that the image of $\tilde{\Delta}_{r}-\lambda I$ is closed, consider a sequence $g_{n}:=(-\tilde{\Delta}_{r}-\lambda I)(f_{n})$ such that $g_{n}\rightarrow g$.
 Similarly to before, we can conclude that since $g_{n}$ converges, then $f_{n}$ converges to some $f$, and thus $g=(-\tilde{\Delta}_{r}-\lambda I)(f)$.
@@
 Corollary 3.13.
-Let $(X,\mathsf{d},\mu)$ be a metric measure space such that for some $r_{0}>0$ both ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) hold for any $r\in(0,r_{0})$. Then
+Let $(X,\mathsf{d},\mu)$ be a metric measure space such that for some $r_{0}>0$ the assumption ($\mathrm{II}_{r}$) holds for any $r\in(0,r_{0})$. Then
 $\lim\limits_{r\downarrow 0}\big{(}\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})\big{)}=+\infty.$
@@
 Proposition 3.14.
-Let $(X,\mathsf{d},\mu)$ be a connected metric measure space satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$.
+Let $(X,\mathsf{d},\mu)$ be a connected metric measure space satisfying ($\mathrm{II}_{r}$) for some fixed $r>0$.
 Then the kernel of $\tilde{\Delta}_{r}$ contains constant functions only, and $\tilde{E}_{r}$ defines a scalar product on
@@
 Consider $f\in L^{2}(X,\mu)\backslash\{0\}$ such that $\tilde{\Delta}_{r}f=0$. Then $\tilde{E}_{r}(f)=0$. This implies that for $\mu$-a.e. $x\in X$,
-[…]
+$\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\left(\frac{f(x)-f(y)}{r}\right)^{2}d\mu(y)=0$ (43)
 which implies, in turn,
 $\mu\left(\{y\in B_{r}(x):f(y)=f(x)\}\right)=\mu(B_{r}(x)).$ (44)
-[…]
+Consider $F\coloneqq\{x\in X:\text{$f(x)$ is a well-defined real number}\}$ and
 $A:=\{x\in F:\mu\left(\{y\in B_{r}(x):f(y)=f(x)\}\right)=\mu(B_{r}(x))\}.$ (45)
 Then
-$\mu(X\backslash A)=0,$ (46)
-Since the measure of any open ball is positive, $X\backslash A$ cannot contain such a ball. This implies that $A$ is a dense subset of $X$. Take $z\in A$ and let $c=f(z)$. Consider
+[…]
+Take $z\in A$ and let $c=f(z)$. Consider
 $I=\{x\in A:f(x)=c\},\quad\quad I^{\prime}=\{x\in A:f(x)\neq c\},$ (47)
@@
 $W:=\bigcup_{x\in I}B_{r}(x),\quad\quad V:=\bigcup_{y\in I^{\prime}}B_{r}(y).$ (48)
-Since $V$ and $W$ are open sets whose union contains $A$ which is dense in $X$, we must have
-$W\cup V=X.$ (49)
-Since $W$ and $V$ form an open cover of $X$, and $X$ is connected, if both $V$ and $W$ are different from the empty set, then there exist $x\in I$ and $y\in I^{\prime}$ such that
+Since $V$ and $W$ are open sets whose union contains $A$ which has full measure in $X$, we must have
+$W\cup V=X,$ (49)
+[…]
 $B_{r}(x)\cap B_{r}(y)\neq\emptyset.$ (50)
@@
 $w\in X$.
-Then $-\tilde{\Delta}_{r}$ has a non-trivial kernel consisting of the constant functions only.
-Moreover, since $-\tilde{\Delta}_{r}$ is non-negative (Remark 3.5), we get the desired property on (42).
-We are now in a position to prove Theorem 1. We recall that the context of this statement is a compact uniformly locally doubling metric measure space $(X,\mathsf{d},\mu)$.
+Then $\tilde{\Delta}_{r}$ has a non-trivial kernel consisting of the constant functions only.
+Moreover, since $\tilde{\Delta}_{r}$ is non-negative (Remark 3.8), we get the desired property on (42).
+∎
+We are now in a position to prove Theorem 1.
+We recall that the context of this statement is a compact uniformly locally doubling metric measure space $(X,\mathsf{d},\mu)$.
+The compactness of the space ensures that ($\mathrm{II}_{r}$) holds for any $r>0$, see Lemma 2.6.
 Proof.
@@
 $\tilde{E}_{r}(f)=\sum_{i=0}^{k}a_{i}^{2}\tilde{E}_{r}(f_{i})\leq\max_{0\leq i\leq k}\tilde{E}_{r}(f_{i}).$
-Let $r_{0},C$ be the parameters of the local doubling property of $(X,\mathsf{d},\mu)$, see Definition 2.1. Then for any $i\in\{0,\ldots,k\}$ and $r\in(0,r_{0})$,
+Let $r_{0},C$ be the parameters of the uniform local doubling property of $(X,\mathsf{d},\mu)$, see Definition 2.8. Then for any $i\in\{0,\ldots,k\}$ and $r\in(0,r_{0})$,
 $\tilde{E}_{r}(f_{i})$ $\leq\frac{\operatorname{Lip}^{2}(f_{i})}{4}\int_{X}\fint_{B_{r}(x)}\underbrace{\left(1+\frac{V(x,r)}{V(y,r)}\right)}_{\leq 1+C^{2}}\underbrace{\frac{\mathsf{d}^{2}(x,y)}{r^{2}}}_{\leq 1}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
@@
 By Corollary 3.13, there exists $r_{k}\in(0,\tilde{r}_{k})$ such that $\tilde{C}<\min\sigma_{\mathrm{ess}}(-\tilde{\Delta}_{r})$ for any $r\in(0,r_{k})$.
-Then Remark 3.6 implies that for such an $r$ the operator $-\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues $\lambda_{0}(-\tilde{\Delta}_{r})\leq\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$ such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$.
+Then Remark 3.11 implies that for such an $r$ the operator $\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues $\lambda_{0}(-\tilde{\Delta}_{r})\leq\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$ such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$.
 That $\lambda_{0}(-\tilde{\Delta}_{r})=0$ follows from Proposition 3.14.
@@
 $\tilde{\lambda}_{1,r}=\min_{f\in\Pi(X,\mu)}\frac{\tilde{E}_{r}(f)}{\|f\|_{2}}$
-where $\Pi(X,\mu)$ is as in (42). Since $-\Delta_{r}$ has a kernel which is $L^{2}$-orthogonal to $\Pi(X,\mu)$ (Proposition 3.9), we have $\tilde{E}_{r}(f)>0$ for any $f\in\Pi(X,\mu)$, hence we get
+where $\Pi(X,\mu)$ is as in (42). Since $\Delta_{r}$ has a kernel which is $L^{2}$-orthogonal to $\Pi(X,\mu)$ (Proposition 3.14), we have $\tilde{E}_{r}(f)>0$ for any $f\in\Pi(X,\mu)$, hence we get
 $\tilde{\lambda}_{1,r}>0.$
@@
 We begin with a result where we use the normalized sinc function, namely
-$\operatorname{sinc}(\rho)\coloneqq\begin{cases}\displaystyle\frac{\sin(\pi\rho)}{\pi\rho}\qquad&\rho\in\mathbb{R}\backslash\{0\},\\ 1&\rho=0,\end{cases}$
+[…]
 and the following notation: for any $p=(p_{1},...,p_{m})\in\mathbb{Z}^{m}$,
@@
 Proof.
-We start by pointing out that for any $r>0$ and $p\in\mathbb{Z}^{m}\backslash\{0\}$,
+[…]
 $\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\neq 1,$
@@
 $\inf_{0\neq p\in\mathbb{Z}^{m}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|=\min_{\begin{subarray}{c}0\neq p\in\mathbb{Z}^{m}\\ |p|_{\infty}<R\end{subarray}}\left|\frac{1}{r^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r)\right)\right|>0.$
-If (56) were to fail, due to the previous line, there would exist sequences $(r_{n})\subset(0,+\infty)$ and $(p^{(n)})\subset\mathbb{Z}^{m}\backslash\{0\}$ such that $r_{n}\to 0$ and
+[…]
 $\lim_{n}\left|\frac{1}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p^{(n)}_{i}r)\right)\right|=0.$ (58)
@@
 For any $x\in\mathbb{T}^{m}$ and $r>0$ small enough, the ball $\tilde{Q}_{r}(x)\subset\mathbb{T}^{m}$ is given by
-$\tilde{Q}_{r}(x)=\pi(Q_{r}(\tilde{x}))$ (67)
-where $\tilde{x}$ is any element in $\pi^{-1}(x)$, and $Q_{r}(\tilde{x})$ is as in (65). Then
+$)\tilde{Q}_{r}(x)=\pi(Q_{r}(\tilde{x}$)) (67)
+[…]
 $\mathbb{L}^{m}(\tilde{Q}_{r}(x))=\frac{\mathcal{L}^{m}(Q_{r}(\tilde{x}))}{2^{m}}=r^{m}$ (68)
@@
 Assume that there exist $r_{n}\rightarrow 0$ and $\{f_{n}\}\subset L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$ satisfying $\int_{\mathbb{T}^{m}}f_{n}\mathop{}\!\mathrm{d}\mathbb{L}^{m}=0$ and $\|f_{n}\|_{L^{2}(\mathbb{T}^{m})}=1$, such that
-$-\tilde{\Delta}_{r_{n}}f_{n}=\lambda_{1}(-\tilde{\Delta}_{r_{n}})f_{n},\quad\quad\lambda_{1}(-\tilde{\Delta}_{r_{n}})\rightarrow 0.$ (70)
+$\tilde{\Delta}_{r_{n}}f_{n}=\lambda_{1}(-\tilde{\Delta}_{r_{n}})f_{n},\quad\quad\lambda_{1}(-\tilde{\Delta}_{r_{n}})\rightarrow 0.$ (70)
 With no loss of generality, we assume that each $r_{n}$ is small enough to ensure that $\tilde{E}_{r_{n},\mathfrak{T}^{m}}$ writes as in (69).
@@
 $=\frac{1}{r_{n}^{2}}\left(f_{n}(x)-\frac{1}{r_{n}^{m}}\bigg{(}1_{\tilde{Q}_{r_{n}}(0)}*f_{n}\bigg{)}(x)\right).$ (71)
-We consider the Fourier decomposition of $f_{n}$, $1_{\tilde{Q}_{r_{n}}(0)}$, and $-\tilde{\Delta}_{r_{n}}f_{n}$, namely
+We consider the Fourier decomposition of $f_{n}$, $1_{\tilde{Q}_{r_{n}}(0)}$, and $\tilde{\Delta}_{r_{n}}f_{n}$, namely
 $f_{n}=\sum_{p\in\mathbb{Z}^{m}}a_{p,n}e_{p}$
 $1_{\tilde{Q}_{r_{n}}(0)}=\sum_{p\in\mathbb{Z}^{m}}b_{p,n}e_{p}$
-$-\tilde{\Delta}_{r_{n}}f_{n}=\sum_{p\in\mathbb{Z}^{m}}c_{p,n}e_{p}$
+$\tilde{\Delta}_{r_{n}}f_{n}=\sum_{p\in\mathbb{Z}^{m}}c_{p,n}e_{p}$
 where $\{e_{p}\}_{p\in\mathbb{Z}^{m}}$ is the orthonormal basis of $L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})$ given by
@@
 $c_{p,n}=\frac{a_{p,n}}{r_{n}^{2}}\left(1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n})\right).$ (74)
-Using Lemma 4.1, for $\0p\in\mathbb{Z}\backslash\{0\}$ we conclude that there exists $\alpha>0$ such that
+[…]
 $\frac{1}{r_{n}^{2}}\left|1-\prod_{i\in J(p)}\operatorname{sinc}(p_{i}r_{n})\right|\geq\alpha.$ (75)
@@
 By Parseval’s identity, and since $f_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$ we have $a_{0,n}=0$,
-$\|-\tilde{\Delta}_{r_{n}}f_{n}\|_{L^{2}(\mathbb{T}^{m})}^{2}=\sum_{p\in\mathbb{Z}^{m}}|c_{p,n}|^{2}\geq\sum_{p\in\mathbb{Z}^{m}\backslash\{0\}}|a_{p,n}|^{2}\alpha^{2}=\alpha^{2}\|f_{n}\|_{L^{2}(\mathbb{T}^{m})}^{2},$ (77)
+[…]
 in contradiction with $\|-\tilde{\Delta}_{r_{n}}f_{n}\|_{L^{2}(\mathbb{T}^{m})}\rightarrow 0$ provided by (70). ∎
@@
 $\|\tilde{f}_{n}\|_{L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})}=1.$
-Let $C_{1},\ldots,C_{2^{m}}$ denote the $2^{m}$ sets of the form $I_{1}\times\ldots\times I_{m}$ where each $I_{i}$ is either $[-1,0]$ or $[0,1]$. For any $j\in\{1,\ldots,2^{m}\}$, set
+Let $C_{1},\ldots,C_{2^{m}}$ denote the $2^{m}$ sets of the form $I_{1}\times\ldots\times I_{m}$ where each $I_{i}$ is either [1,0][-1,0] or $[0,1]$. For any $j\in\{1,\ldots,2^{m}\}$, set
 $\begin{array}[]{cccccc}N_{j}&:&[0,1]^{m}&\rightarrow&C_{j}\\ &&(\xi_{1},\ldots,\xi_{m})&\mapsto&(\varepsilon_{i}\xi_{1},\ldots,\varepsilon_{m}\xi_{m})\end{array}$
-where $\varepsilon_{i}$ is $1$ if $I_{i}=[0,1]$ and $-1$ otherwise.
-Note that $N_{j}$ is an isometry which preserves the Lebesgue measure, and that $T\circ\pi\circ N_{j}$ is equal to the identity.
-Then
+where $\varepsilon_{i}$ is $1$ if $I_{i}=[0,1]$ and 1-1 otherwise. Note that $N_{j}$ is an isometry which preserves the Lebesgue measure, and that $T\circ\pi\circ N_{j}$ is equal to the identity. Then
 $\|\tilde{f}_{n}\|_{L^{2}(\mathbb{T}^{m},\mathbb{L}^{m})}^{2}=\int_{\mathbb{T}^{m}}(f_{n}\circ T)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}$ $=\frac{1}{2^{m}}\int_{[-1,1]^{m}}(f_{n}\circ T\circ\pi)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}$
@@
 Since $\tilde{f}_{n}\in\Pi(\mathbb{T}^{m},\mathbb{L}^{m})$, the latter provides a contradiction with Proposition 4.2, namely
-$0<\lambda_{1}(-\tilde{\Delta}_{\bar{r}_{n},\mathfrak{T}^{m}})\leq 2^{m}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\to 0$
+$0<\lambda_{1}(-\tilde{\Delta}_{\bar{r}_{n},\mathfrak{T}^{m}})\leq 2^{m}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\to 0.$
 Given $x\in\mathbb{T}^{m}$, define
@@
 ${G}_{n}(x)$ $\coloneqq\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{1}{V(x,\overline{r}_{n})}+\frac{1}{V(y,\overline{r}_{n})}\right)$
-×$\qquad\qquad\quad\times\left(\frac{f_{n}(x)-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$
+$\qquad\qquad\quad\times\left(\frac{f_{n}(x)-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$
 where
@@
 If $\pi_{i}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the projection in the $i$-th coordinate, we conclude
-$\pi_{i}(R_{n,k}(x))=\begin{cases}[0,\overline{r}_{n}+|\overline{x}_{i}|]&\text{ if }(1)\\ [-\overline{r}_{n}+|\overline{x}_{i}|,1]&\text{ if }(2)\\ [-\overline{r}_{n}+|\overline{x}_{i}|,1]&\text{ if }(3)\\ [-\overline{r}_{n},\overline{r}_{n}]&\text{ if }(4)\end{cases}$ (87)
-$\pi_{l}(R_{n,k}(x))=\begin{cases}\pi_{i}(R_{n,l}(x))=\pi_{i}(R_{n,k}(x))&\text{ if }\pi_{i}(C_{j})=\pi_{i}(C_{k})\\ [0,\overline{r}_{n}-|\overline{x}_{i}|]&\text{ if }(1)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ [2-\overline{r}_{n}-|\overline{x}_{i}|,1]&\text{ if }(2)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ [2-\overline{r}_{n}-|\overline{x}_{i}|,1]&\text{ if }(3)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ \emptyset&\text{ if }(4)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right).\end{cases}$ (88)
+[…]
+[…]
 Thus we conclude that for all $i\in\{1,...,m\}$ and $l\in\{1,...,2^{m}\}$ we have $\pi_{i}(R_{n,l}(x))\subset\pi_{i}(R_{n,k}(x))$, and since these sets are rectangles, we conclude equation (86).
@@
 $\leq\int_{\tilde{Q}_{\overline{r}_{n}}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)$ (90)
-×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (91)
+$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (91)
 $=\sum_{j=1}^{2^{m}}\int_{\tilde{Q}_{\overline{r}_{n}}(x)\cap\pi(C_{j})}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)$ (92)
-×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (93)
+$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (93)
 $=\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j}}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(\pi(y)),\overline{r}_{n})}\right)$ (94)
-×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(y)))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (95)
+$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(y)))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (95)
 For each integral, change coordinates by $N_{j}$ to conclude
 $\tilde{G}_{n}(x)$ $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{N_{j}^{-1}(\pi^{-1}(\tilde{Q}_{\overline{r}_{n}}(x))\cap C_{j})}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(\pi(N_{j}(y))),\overline{r}_{n})}\right)\times$ (96)
-×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(N_{j}(y))))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y).$ (97)
+$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(N_{j}(y))))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y).$ (97)
 We have that
@@
 $\tilde{G}_{n}(x)$ $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{R_{n,j}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(T(y),\overline{r}_{n})}\right)\times$ (99)
-×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (100)
+$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (100)
 $\leq\frac{1}{2^{m}}\sum_{j=1}^{2^{m}}\int_{R_{n,k}(x)}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(y,\overline{r}_{n})}\right)\times$ (101)
-×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (102)
+$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (102)
 $=\int_{[0,1]^{m}\cap Q_{\overline{r}_{n}}(T(x))}\left(\frac{2^{m}}{V(T(x),\overline{r}_{n})}+\frac{2^{m}}{V(y,\overline{r}_{n})}\right)\times$ (103)
-×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (104)
+$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (104)
 $=2^{m}G_{n}(T(x)).$ (105)
@@
 In this smooth context, the function $V(\cdot,r)$ is obviously continuous for any $r>0$.
-Since $M$ is compact, this implies that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$).
+Since $M$ is compact, this implies that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies ($\mathrm{II}_{r}$) and ($\mathrm{I}_{r}$).
 Then Lemma 3.3 applies and ensures that $\tilde{\Delta}_{r}$ is a bounded self-adjoint operator acting on $L^{2}(M,\mathrm{vol}_{g})$.
@@
 5.1. Convergence in the sense of distributions
-Recall that $C_{m}$ is defined in (3).
-For any $x\in M\backslash\partial M$, we let $\exp_{x}$ be the exponential map centered at $x$.
-We identify $T_{x}M$ with $\mathbb{R}^{m}$ and write $\mathbb{B}^{m}_{r}(v)$ for the Euclidean ball in $\mathbb{R}^{m}$ centered at $v$ with radius $r>0$.
-Then there exists $\delta>0$ such that the restriction of $\exp_{x}$ to $\mathbb{B}^{m}_{\delta}(0)$ is a diffeomorphism onto its image; recall that the injectivity radius $\mathrm{i}_{M}(x)$ of $M$ at $x$ is the supremum of the set of such numbers $\delta$.
-We let $J_{x}$ be the Radon-Nikodym derivative of the measure $(\exp_{x}^{-1})_{\#}\mathrm{vol}_{g}$ with respect to the Lebesgue measure $\mathcal{L}^{m}$.
-It is well-known that for any $\xi\in\mathbb{B}_{\mathrm{i}_{M}}^{m}(0)$,
+[…]
 $J_{x}(\xi)=1+\underline{O}_{K}(|\xi|^{2})$
@@
 5.3. Uniform bound
-We wish now to provide a uniform $L^{\infty}$ bound for the functions $-\tilde{\Delta}_{r}\psi$ where $r$ is in a neighborhood of zero.
+We wish now to provide a uniform $L^{\infty}$ bound for the functions $\tilde{\Delta}_{r}\psi$ where $r$ is in a neighborhood of zero.
 Let us first consider the case $\partial M=\emptyset$. From Proposition 5.3, we have uniform convergence
@@
 $(\psi^{*}g_{t})_{\xi}(v,w)=v^{T}\cdot A_{(\xi,t)}\cdot w,$ (122)
-for some positive definite, symmetric $(m-1)$-square matrix $A_{(\xi,t)}$.
+for some positive definite, symmetric (m1)(m-1)-square matrix $A_{(\xi,t)}$.
 From the non-degeneracy of the metric and a Lipschitz bound, we have that there exist $C,\tilde{c}>0$ such that for all $t,s\in[0,\epsilon],\xi\in[-3,3]^{m-1},v\in\mathbb{R}^{m}$ we have
@@
 Proof.
-Consider $\tilde{c}>0$ given by (123).
-We know that there exists $M>0$ such that the map $\sqrt{\cdot}:[\tilde{c},+\infty)\rightarrow\mathbb{R}$ is $M$-Lipschitz.
-By homogeneity in $|v|$ of (125), we can assume that $|v|=1$.
-Then the Lipschitz condition and (124) yield
+[…]
 $\left|O(\xi,t,s,v)\right|$ $=\left|\left[(\psi^{*}g_{t})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)-\left[(\psi^{*}g_{s})\oplus d\tau^{2}\right]_{(\xi,t)}^{\frac{1}{2}}(v,v)\right|$ (127)
@@
 $(\Phi^{*}g)_{(\xi,t)}=(\psi^{*}g_{t}\oplus\mathop{}\!\mathrm{d}\tau^{2})_{(\xi,t)}.$ (132)
-Set $\mathbb{H}^{m}\coloneqq\{v\in\mathbb{R}^{m}:v_{m}\geq 0\}$. Let $L(\tilde{\gamma})$ be the Euclidean length of a curve $\tilde{\gamma}:[0,1]\to\mathbb{R}^{m}$.
+[…]
 Claim 2.
@@
 […]
-$=\sqrt{\tilde{c}}\,L(\tilde{\gamma}).$ (139)
+[…]
 Now we claim that:
@@
 $\Psi_{(\xi,t)}(\zeta,s)=\Psi(\xi,t,\zeta,s)$
-to see $\Psi$ as a function of the two last variables only, the two first being frozen. Observe that for any $(\xi,t,\zeta)$ in $[-1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)$,
+to see $\Psi$ as a function of the two last variables only, the two first being frozen. Observe that for any $(\xi,t,\zeta)$ in $1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)$,
 $(g_{t})^{\frac{1}{2}}_{\psi(\xi)}((d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta,(d\psi)_{\xi}A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ $=(\psi^{*}g_{t})_{\xi}^{\frac{1}{2}}(A_{(\xi,t)}^{-\frac{1}{2}}\zeta,A_{(\xi,t)}^{-\frac{1}{2}}\zeta)$ (164)
@@
 $B_{r_{0}}(\Psi_{(\xi,t)}(0,t))=B_{r_{0}}(\Phi(\xi,t))\subset\Psi_{(\xi,t)}(\mathbb{B}^{m-1}_{\delta}(0)\times[0,\epsilon]).$ (175)
-By compactness of $[-2,2]^{m-1}\times[0,\epsilon/2]$ and continuity of the maps $\Psi_{(\xi,t)}^{-1}$, we get that there exists a common $r_{0}>0$ such that the previous holds for any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$.
+By compactness of $2,2]^{m-1}\times[0,\epsilon/2]$ and continuity of the maps $\Psi_{(\xi,t)}^{-1}$, we get that there exists a common $r_{0}>0$ such that the previous holds for any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$.
 Consider $r<r_{0}$ and $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, then
@@
 $G_{r}(x)=\frac{1}{r^{2}}\int_{\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))}\frac{1}{\mathrm{vol}_{g}(B_{r}(\Psi_{(\xi,t)}(\zeta,s)))}$ (208)
-×$\times\left(f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))\right)\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)$ (209)
+$\times\left(f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))\right)\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)$ (209)
 $\begin{split}G_{r}(x)&=\frac{1}{r^{2}}\int_{\Psi_{(\xi,t)}^{-1}(B_{r}(\Psi_{(\xi,t)}(0,t)))}\frac{f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\times\\ &\quad\times\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)+O(1)\end{split}$ (by 7 and 8)
@@
 $\frac{1}{r^{2}}\int_{\mathbb{B}^{m}_{r}(0,t)\cap\mathbb{H}^{m}}\frac{\partial_{m}(f\circ\Psi_{(x,t)})(0,t)(s-t)}{\mathcal{L}^{m}(\mathbb{B}^{m}_{r}(0,s)\cap\mathbb{H}^{m})}\ \mathop{}\!\mathrm{d}\mathcal{L}^{m-1}(y)\mathop{}\!\mathrm{d}\mathcal{L}(s)=O(1),$ (214)
-which shows that $G_{r}(x)=O(1)$ for $x$ such that $d(x,\partial M)<2r$.
+which shows that $G_{r}(x)=O(1)$ for $x$ such that $d(x,\partial M)<2r$. ∎
 6. Spectral convergence
@@
 $=\iint_{\mathcal{Q}^{2}}\tilde{a}_{r,\mathfrak{M}}(\Phi(x),\Phi(y))\left(\frac{f(\Phi(x))-f(\Phi(y))}{r}\right)^{2}$
-×$\qquad\qquad\qquad\qquad\qquad\qquad\times\sqrt{\det(g_{x})\det(g_{y})}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
+$\qquad\qquad\qquad\qquad\qquad\qquad\times\sqrt{\det(g_{x})\det(g_{y})}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
 $\geq\iint_{\mathcal{Q}^{2}}\frac{c}{C}\tilde{a}_{cr,\mathfrak{Q}}(x,y)\left(\frac{f(\Phi(x))-f(\Phi(y))}{r}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
@@
 $\partial M\neq\emptyset$ ), $k$ is a positive integer, $\mu_{k}$ is the $k$-th lowest Laplace (resp.
-Neumann) eigenvalue of $\Delta_{g}$, and $f_{k,r_{n}}$ is an eigenfunction of $-\tilde{\Delta}_{r_{n}}$ associated with the $k$-th eigenvalue $\lambda_{k}(-\tilde{\Delta}_{r_{n}})$ of this operator.
+Neumann) eigenvalue of $\Delta_{g}$, and $f_{k,r_{n}}$ is an eigenfunction of $\tilde{\Delta}_{r_{n}}$ associated with the $k$-th eigenvalue $\lambda_{k}(-\tilde{\Delta}_{r_{n}})$ of this operator.
 Proof.